### v3

# DiffArtist: Towards Structure and Appearance Controllable Image Stylization

Abstract.
Artistic style includes both structural and appearance elements. Existing neural stylization techniques primarily focus on transferring appearance features such as color and texture, often neglecting the equally crucial aspect of structural stylization. In this paper, we present a comprehensive study on the simultaneous stylization of structure and appearance of 2D images. Specifically, we introduce DiffArtist, which, to the best of our knowledge, is the first stylization method to allow for dual controllability over structure and appearance. Our key insight is to represent structure and appearance as separate diffusion processes to achieve complete disentanglement without requiring any training, thereby endowing users with unprecedented controllability for both components. The evaluation of stylization of both appearance and structure, however, remains challenging as it necessitates semantic understanding. To this end, we further propose a Multimodal LLM-based style evaluator, which better aligns with human preferences than metrics lacking semantic understanding. With this powerful evaluator, we conduct extensive analysis, demonstrating that DiffArtist achieves superior style fidelity, editability, and structure-appearance disentanglement. These merits make DiffArtist a highly versatile solution for creative applications. Project homepage: https://github.com/songrise/Artist.
Generative art, Text-driven stylization, Structure and appearance, Stylization evaluation, Multimodal LLM applications,
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Arxiv; 2025; NYisbn: 978-1-4503-XXXX-X/2018/06submissionid: ccs: Computing methodologies Image processingccs: Applied computing Fine artsccs: Computing methodologies Artificial intelligence
1. Introduction
Figure 1. DiffArtist enables disentangled and fine-grained control of style strength from two orthogonal perspectives: structure and appearance. The prompt is “The Dream, by Picasso.”
The essence of an artistic style lies not only in its visual appearance but also the structural elements (Gombrich and Gombrich, 1995; Jung et al., 2024). For example, the fragmentation of figures in Picasso’s Cubist works and the undulating sky in Van Gogh’s “Starry Night”, each contributing distinctly to their artistic expression. To automate the generation of artworks, existing neural stylization approaches (Huang and Belongie, 2017; Deng et al., 2022; Kwon and Ye, 2022; Wang et al., 2022; Huang et al., 2024; Cui et al., 2023) have predominantly focused on manipulating appearance-level attributes such as color and textures. The structural elements in the source image, however, are often modeled as the “content” and are explicitly preserved (Wang et al., 2024b; Wen et al., 2023; Yang et al., 2022). Consequently, these methods typically lack the capacity to curate meaningful structural-styles, substantially limiting their artistic potential and customizability.
Unlike appearance, structural stylization is inherently much more challenging as it demands artistically meaningful variations that (1) are aligned with the target style, (2) harmoniously integrated with the original image, and (3) without violating the semantic integrity of the source content. These objectives are at a higher, potentially semantics-level than the appearance-style, and, therefore, are less readily achieved. While descriptors like AdaIN (Huang and Belongie, 2017) and Gram loss (Gatys et al., 2015) may suffice for appearance-style modeling, the lack of adequate structure representation and a structure-style evaluator represent major obstacles in the development of structural stylization techniques. This challenge becomes even more pronounced in contemporary multimodal generation scenarios, where the style as textual prompt lacks a direct visual counterpart (Huang et al., 2024; Wang et al., 2023a; Cui et al., 2023; Hertz et al., 2024).
In search of the dual controllability of structure and appearance elements, the recent development of image generation models, in particular the diffusion models (DMs) has provided a promising solution. Compared with traditional stylization, the generative process allows for the sampling of stylized images of greater diversity in both structure and appearance. Building upon DMs, the stylization problem can re-framed as controlled image generation, where the reference image and style are additional conditions for sampling of stylized images. However, achieving this control in an artistically meaningful way is far from straightforward, as the diffusion process inherently entangles the generation of structure and appearance, as analyzed in this paper. This entanglement presents a major challenge for full controllability, resulting in a tradeoff between structural and appearance style strength (i.e., the S-A tradeoff). Consequently, existing diffusion-based approaches are either prone to undesired content modification (Chung et al., 2024; Lyu et al., 2023; Wang et al., 2024b) or suffer from constrained appearance-style strength (Wang et al., 2023b). Achieving dual controllability in the stylization thus remains an open question.
In this paper, we present the first comprehensive study for disentangled structure and appearance control for 2D stylization. We first present an styler, DiffArtist, which, to the best of our knowledge, is the first to enable the dual controllability for text-driven stylization. DiffArtist is based on explicit disentanglement of the structural and appearance generation in the diffusion process. It is plug-in with U-Net-based DMs, requiring neither optimization nor additional adapters like ControlNet (Zhang et al., 2023a) or IP-Adapter (Ye et al., 2023) (i.e., zero-shot), making it highly efficient for user personalization. This superior controllability overcomes the S-A tradeoff in DMs. Compared with conventional methods and other DM-based methods (e.g, (Wang et al., 2024a; Lin et al., 2024; Zhang et al., 2023a)), we demonstrate DiffArtist achieves significantly better dual controllability without deterioration in semantics, as exemplified in Fig. 1. Aside from structural stylization, it also allows appearance stylization with un-represented high fidelity, when configured with strict structure preservation, thereby unleashing full customization power.
In addition to the proposed styler, we further explore the development of art evaluators with human-aligned perceptions of structure and appearance style. The challenges for this exploration are threefold: (1) Structural style components are at a higher level than appearance and cannot be evaluated in isolation without considering their integration with content. (2) Evaluators must possess contextual awareness to detect structural artifacts that compromise semantic coherence. (3) Evaluators should also effectively associate textual style instructions with corresponding visual features for cross-modal evaluation. Multimodal LLMs (MLLMs), known for their robust zero-shot semantic-level understanding capabilities, emerge as an optimal solution to address these challenges. To this end, we propose (to the best of our knowledge) the first MLLM-based metrics for evaluating the structure and appearance in artistic stylization. We empirically demonstrate that this metric aligns significantly better with human artistic perception than existing metrics (Kirstain et al., 2023; Wang et al., 2004; Radford et al., 2021), paving the way for more human-aligned style evaluation. We summarize our contributions as follows:
  1. (1)
    We propose DiffArtist, the first stylization method that enables the dual controllability of structure and appearance-related style.
  2. (2)
    We present a novel MLLM-based evaluator for evaluation of structure and appearance in artistic stylization, which aligns better with human perception.
  3. (3)
    Extensive experiments demonstrate DiffArtist achieves superior stylization fidelity, with highly editable and disentangled control.
2. Related Works
The Stylization of Structure and/or Appearance. Structure and appearance collectively define the style of a visual representation. Existing neural stylization methods (Huang and Belongie, 2017; Deng et al., 2022; Kwon and Ye, 2022; Wang et al., 2022; Huang et al., 2024; Cui et al., 2023) has predominately focused on appearance stylization, generally achieved via encoder-decoder architecture. Only a few papers (Kim et al., 2020; Zhou et al., 2024) focus on transferring structural style components between 2D images. This can be achieved via image deformation. For instance, one may calculate a correspondence between the structure of content and style image, to perform a non-rigid deformation. However, these methods usually operate images in specific domains like portrait, requiring a stylized image as reference. They also do not allow for dual controllability. The dual controllability has recently been explored in image editing literature (Lin et al., 2024), yet they focus on targeted editing of realistic photos at single control level, instead of harmonious transfer between photo and artworks with high editablity. The structure stylization is more widely explored in 3D stylization, which has an explicit formulation of shapes (e.g, vertices (Michel et al., 2022), density (Jung et al., 2024)) and appearance (e.g., texture (Chen et al., 2023b) map, radiance (Wang et al., 2023a)), yet they cannot be directly applied to the 2D case. In this paper, we explore the first dual controllable 2D image stylization method.
Text-Driven Image Stylization. Text-driven image stylization aims to stylize a source image according to style prompts. Early methods achieve it by optimizing certain image representation (Michel et al., 2022; Kwon and Ye, 2022; Wang et al., 2023a; Gal et al., 2022; Kim et al., 2022) with a multimodal similarity loss, usually implemented as the CLIP loss (Radford et al., 2021). Recently, it was discovered that text-to-image (T2I) DMs could also be adapted for similar optimization schemes (Poole et al., 2022; Hertz et al., 2023; Jiang et al., 2023; Kawar et al., 2023). These optimization-based methods are costly and slow, motivating recent methods explore the feed-forward paradigm. Instruct-Pix2Pix (Brooks et al., 2023) tune the diffusion model with a language model for generalized editing tasks. Diffstyler (Huang et al., 2024) learn a content and style-specific denoiser for disentanglement. FreeStyle (He et al., 2024) modulate U-Net feature for training-free stylization. Concurrent with this exploration is the stylized image generation. While related, they focus a different setting where the style is extracted from an image and the content is a prompt (Hertz et al., 2024; Wang et al., 2024a, b; Gao et al., 2024; Chen et al., 2023a). In this work, we focus on the structure and appearance control in text-driven stylization scenarios.
Quantitative Evaluation of Style Transfer. Traditional metrics such as Gram Loss (Gatys et al., 2016), LPIPS (Zhang et al., 2018), and Fréchet Inception Distance (FID) (Heusel et al., 2017) evaluate quality of stylization from various perspectives. These metrics are at a relatively low-level, and are found to be misaligned with the holistic human artistic perception (Bo et al., 2018; So, 2023; Ioannou and Maddock, 2024; Chen et al., 2024; Yi et al., 2023; Wright and Ommer, 2022). To build art-specific evaluators, recent methods like ArtFID (Wright and Ommer, 2022) and ArtScore (Chen et al., 2024) attempt to quantify ”artness” of images, yet these methods cannot handle the open-vocabulary text-driven stylization, and they do not evaluate structure and appearance separately. In this paper, we propose a semantic-level MLLM-based evaluator to assess the structure and appearance fidelity, which aligns better with human perception.
3. Methodology
Figure 2. Overview of DiffArtist. We disentangle the structure and appearance of image as separate diffusion processes. All branches share the same U-Net parameters, and perform full denoising of $\tau$ steps. The feature pyramid from structure delegation is injected into the main branch for structure control, and the semantic-aware appearance is controlled via AdaIN.
3.1. Objective: Dual Controllability
Given a content image $I$, a style prompt $y$, and a U-Net-based DM $\mathcal{G}$, our objective is to control $\mathcal{G}$ to derive a distribution of stylized images $\hat{I}\sim\mathcal{G}(I,y)$ that harmoniously compose the provided content and style. Specifically, we aim to achieve disentangled controllability in both structure and appearance-related style. Our definition of structure and appearance in 2D imagery is mainly based on fine-art (Gombrich and Gombrich, 1995). A formal definition of them relates to visual semiotics (Van Leeuwen and Jewitt, 2000) and extend beyond the scope of this paper. Generally speaking, structure corresponds to the geometry, like contours and curvatures, while appearance corresponds to patterns like strokes and color pallette. We also aim to develop an evaluator $\mathcal{E}$, which can evaluate the fidelity of structure and appearance style in a way aligned with human perception.
The remaining parts of this section are organized as below. In Sec. 3.2, we review the basics of inversion-based image manipulation. In Sec. 3.33.4 we explain the motivation and design of control at a high-level, and details are described in Sec. 3.5. Sec. 3.6 outlines the proposed MLLM-based structure and appearance evaluators.
3.2. Preliminary: DDIM Inversion
To stylize a real image $\mathbf{x}_{0}:=I$ using DMs, inversion-based methods first approximate the noise latent $\mathbf{\tilde{x}}_{1:T}$ of it, typically using techniques such as DDIM inversion (Song et al., 2020). With the inverted latents, stylization can be achieved through generation with altered conditions (usually given as a style prompt $y$). Specifically, one may start with an intermediate step $\mathbf{x}_{\tau}$ (i.e., control point), where $\tau\in[1,T]$ for iterative DDIM sampling. Each denoising step is defined as:
$\mathbf{x}_{t-1}$ $=\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t;y)}{\sqrt{\alpha_{t}}}\right)$
(1) $\qquad+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(\mathbf{x}_{t},t;y),$
where $t$ is the timestep, and $\alpha_{1:T}$ is a specific noise schedule. The assumption in this paradigm is that with a proper $\tau$, the resulting stylized image $\hat{I}:=\hat{\mathbf{x}}_{0}$ harmoniously integrate the content of original image ${\mathbf{x}_{0}}$ with the style in prompt $y$.
3.3. Structure and Appearance in Noise Space
Existing neural stylization methods generally view an image as a composition of certain “content” and “style” components  (Li et al., 2017; Huang and Belongie, 2017; Wang et al., 2023b; Kotovenko et al., 2019; Sanakoyeu et al., 2018; Zhang et al., 2023b, 2020), where the style usually refers to the low-level visual features such as appearance. As this paper motivates to stylize both structure and appearance, we adopt a different modeling to decompose image as appearance and structure components: $\mathbf{x}_{0}=\mathcal{G}_{0}(\mathbf{z}^{a}_{0},\mathbf{z}^{s}_{0})$, where $\mathcal{G}_{0}=(\cdot,\cdot)$ is a composition function, $\mathbf{z}^{a}_{0}$ and $\mathbf{z}^{s}_{0}$ are the latent appearance and structure factorization, respectively. This is a “static” perspective. In the diffusion process, the distribution of image $x_{0}$ is tied with the intermediate distributions in $\mathbf{x}_{1:T}$, where a denoiser $\epsilon_{\theta}$ learns the transition $q_{\theta}(x_{t-1}|x_{t},y)$ via $\epsilon$-prediction. We therefore posit similar factorization of noise-residuals $\epsilon_{\theta}$:
(2) $\epsilon_{\theta}(\mathbf{x}_{t},t;y)=\mathcal{G}_{t}\left(\kappa_{t},\psi_{t}\right),$
where $\kappa_{t}$ and $\psi_{t}$ denote the structure and appearance representation at diffusion time-step $t$ (detailed later), and $\mathcal{G}_{t}(\cdot,\cdot)$ is the noise-space composition function at time $t$.
3.4. Structure and Appearance as Delegate Process
Controlling the structure and appearance-level stylization in DMs presents unique challenges due to inherent entanglement. We provide both theoretical and empirical evidence for this phenomenon in Appendix. B, revealing the technical challenge of S-A tradeoff in DM. The key insight from this analysis is that solidary control point $\mathbf{x}_{\tau}$ used in existing DMs-based stylization does not suffice for S-A disentanglement. Consequently, we argue that the structure and appearance shall be represented and controlled over complete and separate diffusion trajectories.
We leverage two additional diffusion processes with shared information for dual controllability, which we call delegate branches. We initialize the structure and main branch from the inverted noise, while appearance delegation starts from a sampled noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. With these delegations, we are able to control the stylization over the entire diffusion process. The controlled main stylization branch can be denoted as:
(3) $\epsilon^{m}_{\theta}(\mathbf{x}_{t},t;y,\kappa^{s},\psi^{a})=\mathcal{G}_{t}\left(\kappa^{s}\circ\kappa^{m},\psi^{a}\diamond\psi^{m}\right),\quad t\in[1,T]$
where the superscripts $s$, $a$, and $m$ denote the factorization extracted from the structure, appearance delegation, or main branch, respectively. The $\circ$ and $\diamond$ are two non-commutative control operators.
3.5. Structure and Appearance Representations in Denoising U-Net
In this section we formulate the $\kappa$ and $\psi$ in U-Net-based denoiser for enabling structure and appearance disentangled style control.
Pyramidal Structure Representation. A multi-scale, and semantic-rich image representation is essential for structure stylization at different level of abstractness. To achieve this, we find the hidden features in the ResBlock of denoising U-Net encode appearance-invariant image semantics across varying $t$ (see Fig. 3). Formally, we denote the hidden feature of ResBlock as $h_{i}(\mathbf{x}_{t})$, where $i\in\{1,2,\ldots,N_{res}\}$ indexes the ResBlocks in the decoder with varying spatial resolution up to $N_{res}$. Stacking the hidden feature from all layers forms a feature pyramid as the structure representation of $\mathbf{x}_{0}$ at $t$:
(4) $\kappa_{t}$ :=$:=\{h_{i}(\mathbf{x_{t}})\}_{i\in S_{res}},$
[…]
and $\text{and }S_{res}$$\subseteq\{1,2,\ldots,N_{res}\}.$
Technically, the representation above captures the semantics of $x_{0}$ at multiple spatial resolutions and (almost) full signal-to-noise ratio (SNR), making it ideal for structural stylization that requires semantic awareness. This is fundamentally different with latent-based, solidary representation (e.g., $\mathbf{x}_{\tau}$) and pixel-level representation (e.g., ControlNet, IP-Adapter), which are uni-resolution, at single SNR level, and semantically-poor. Consequently, as evidenced in our experiment, these methods fail to produce meaningful structure styles and are prone to deterioration of original semantics, both in terms of structure and appearance. We implement the content control operator $\circ$ as injection (i.e., $a\circ b=a$).
Semantic-aware Appearance Representation. Akin to structure, the appearance-related style also must also be harmoniously aligned with the image semantics. However, until now, the appearance delegation denoise from a random noise and hence has no interaction with the source image. To solve this, we propose Structure-to-Appearance injection (S2A) to propagate the high-level semantics into the generation of appearance to avoid spatial misalignment. Specifically, we inject self-attention value $V^{s}$ from early layers of $\epsilon^{s}_{\theta}$ to $\epsilon^{a}_{\theta}$.
We represent the appearance of the target style as self-attention maps from all layers in the $\epsilon^{s}_{\theta}$. Let $N_{attn}$ denote the total number of attention blocks within the U-Net decoder, $S_{s2a}\subseteq\{1,2,\ldots,N_{attn}\}$ be the selected blocks for s2a injection. The style representation is:
(5) $\sigma^{a}_{t}:=\{A^{s}_{i}\}_{i\in[1,2,\cdots,N_{attn}]}$
[…]
Figure 3. ResBlock feature map visualization. We apply t-sne to visualize the feature map of different feature maps in U-Net. The hidden features $h(\mathbf{x})$ better preserves the semantics than the ResNet feature $f(\mathbf{x})$ throughout all $T$.
Figure 4. Qualitative comparison with existing methods. We compare our work with representative text-driven image manipulation method in (a), and image-based stylization methods in (b). Stylized images generated by DiffArtist produce high-fidelity structural and appearance-level style with semantic integrity. We suggest readers for more visualizations in Appendix. A.
Inspired by StyleAligned (Hertz et al., 2024) we design the appearance-style control operator $\diamond$ as the AdaIN (Huang and Belongie, 2017),
(6) $a\diamond b=\operatorname{Var}(b)(\frac{a-\operatorname{Mean}(a)}{\operatorname{Var}(a)})+\operatorname{Var}(b).$
The entire control process is visualized in Fig. 2. Adjusting the control layer $S_{res}$ and style strength in $\epsilon_{t}^{s}$ enables disentangled control for structure and appearance, respectively.
3.6. MLLMs as Stylization Evaluator
Recent research demonstrates a strong semantic-level multimodal understanding of MLLMs (Zhang et al., 2024; Kojima et al., 2022; Zhang et al., 2023c). Inspired by these findings, we leverage state-of-the-art MLLMs as strong zero-shot artistic evaluators. Due to the inherent subjectivity in artistic evaluation, we argue that obtaining an absolute score for structure or appearance fidelity can be challenging. Consequently, we evaluate them in a relative way, where the results from all compared methods are visible to MLLM for ranking. Besides, we aim to measure the fidelity, instead of the more subjective visual-appeal [^1] Note that, visual-appeal is also not equivalent to the concept of aesthetics. (Gombrich and Gombrich, 1995).
[…]
4. Experiments
Table 1. Quantitative comparison against existing methods. We show conventional (in gray font) and MLLM-based metrics for representative methods. For each metric,   indicates the best score,   indicates the second best score , and   indicates the third best score (best viewed in color). Win rate means the percentage that our method wins in pair-wise comparison.
Metric Ours DDIM SD PnP P2P InstructP2P ControlNet InstantStyle DiffStyler CLIPStyler
Inference time (sec) 10.5 9.7 3.9 55.3 29.1 9.2 7.8 7.8 18.2 24.2
Training & adapter free
LPIPS $\downarrow$ 0.52 0.57 0.76 0.67 0.47 0.42 0.65 0.59 0.71 0.46
CLIP Score (Radford et al., 2021) $\uparrow$ 25.91 25.25 27.46 24.89 23.48 21.94 24.93 22.85 25.79 27.14
PickScore (Kirstain et al., 2023) $\uparrow$ 20.51 20.58 20.68 20.34 20.50 20.06 20.46 19.97 19.24 20.13
Structure (MLLM) $\uparrow$ 0.61 0.22 0.29 0.52 0.65 0.60 0.58 0.56 0.35 0.51
Appearance (MLLM) $\uparrow$ 0.67 0.46 0.31 0.60 0.47 0.59 0.55 0.67 0.30 0.59
Avg. (MLLM) $\uparrow$ 0.64 0.34 0.30 0.56 0.56 0.60 0.57 0.62 0.33 0.55
Structure Win (Human) $\uparrow$ - 78.2% 62.4% 64.7% 57.3% 62.2% 71.2% 59.8% 81.3% 73.0%
Appearance Win (Human) $\uparrow$ - 74.2% 86.4% 62.0% 73.7% 68.7% 75.0% 60.1% 85.3% 76.3%
4.1. Experiment Setup
Implementation details. The pretrained Stable Diffusion 2.1  [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1 is used in our main experiment. We perform DDIM sampling with $T=50$ steps. During the inversion, the intermediate noise predictions are recorded, which are used to overwrite the input of denoiser of the content delegation at every steps. Further implementation details can be found in Appendix B.
We default $S_{\text{res}}$ to be the first four ResNet layers ([1, 2, 3, 4]), and $S_{\text{s2a}}$ as the first two attention layer features ([1, 2]), unless otherwise specified. The classifier-free guidance (CFG) scale is set as $7.5$. Note that the default parameter corresponds to only moderate structural variations, used to set a fair comparison with other works to avoid per-image tuning. However, it should be noted that this paper focuses on improved controllability to allow for customization, allowing user to personalize according to their taste. Experiments are conducted with a single RTX 4090-D GPU. The inversion process takes approximately 2 seconds, and the subsequent stylization process takes around 8 seconds.
Compared methods. We compare our method against existing text-driven stylization and manipulation methods: DDIM Inversion (Song et al., 2020), CLIPStyler (optimization-based) (Kwon and Ye, 2022), DiffStyler (Huang et al., 2024), Plug-and-Play (PnP), Prompt2Prompt (with null text inversion) (P2P)  (Mokady et al., 2023; Hertz et al., 2022), ControlNet (Zhang et al., 2023a), and InstructPix2Pix (Brooks et al., 2023). Additionally we consider a baseline named SD, which generate image with the Stable Diffusion according to $y$.
We also indirectly compare our method with reference-based stylization methods, including AdaIN (Huang and Belongie, 2017), ArtFlow (An et al., 2021), StyTr2 (Deng et al., 2022), StyleID (Chung et al., 2024), StyleAligned (Hertz et al., 2024) (with ControlNet) and InstantStyle (Wang et al., 2024a) (with ControlNet). Images generated by SD are used as reference.
Conventional Metrics: LPIPS (Zhang et al., 2018) measures the content preservation by calculating the perceptual distance from source and stylized image. For style fidelity, we leverage CLIP Score (Radford et al., 2021) and Pick Score (Kirstain et al., 2023), both quantify the alignment between the stylized image $\hat{I}$ and prompt $y$. We also include a human study crowd-sourced from $n_{1}=200$ users and report the average preference rate.
MLLM-based Metrics: We rank the stylization fidelity in ascending order, and normalize the rankings across all test samples. Therefore, a score closer to 1 indicate stronger fidelity. We use Gemini-v2.0-flash is used due to its strong multimodal task-solving capability and empirically superior alignment with human. The full prompt templates can be found in Appendix C.
4.2. Comparisons
Qualitative Comparisons. We first provide a comprehensive comparison against previous text and reference-based stylization methods, visualized in Fig. 4-(a, b). (a): When compared with text-driven methods, DiffArtist is the best at following the style prompt while maintaining content integrity. Our method enables harmonious structural variations, such as pixelation, without compromising intricate details like facial identity and hair. In contrast, the compared methods (e.g., DiffStyler, Plug-and-Play) may introduce undesired modifications that violates semantics or produce misaligned styles (e.g., InstructP2P, ControlNet). (b): When broadly compared with reference-based methods, our method still stands out for its capability to curate both appearance and structural variations. To fully demonstrate the results, we suggest readers for additional visualizations in Appendix A.
Quantitative Comparison. For our quantitative evaluation, we first sample 50 art styles from WikiART, with both abstract (e.g., “Cubism”) and realistic styles (e.g, “High-renaissance”), which are diversified by GPT-4o in terms of description. This sets a wide spectrum of styles to align with real-world applications. The content comprises 50 images from MSCOCO (Lin et al., 2014) and 50 photorealistic images generated by another model (Ideogram, 2024). For each content image we randomly draw $10$ style prompts from all possible styles, resulting in a total of $1,000$ combinations for comparison. The comparative results are presented in Tab. 1.
For conventional metrics, DiffArtist achieves an LPIPS of 0.52, a CLIP Score of 25.91, and a PickScore of 20.51, outperforming most of the compared methods. However, these metrics discard either the content or style reference, and they do not explain stylization quality in terms of structure and appearance. For example, the baseline SD has be best CLIP Score and PickScore but it completely discard the content information in the original image. We include these scores solely for reference.
When evaluated with MLLM-based metrics, DiffArtist attains the highest average score. Specifically, our method achieves the second-highest structure score of 0.61, demonstrating DiffArtist’s effectiveness in generating structural styles. The only stronger baseline is P2P, which is designed for image editing, with relatively weak stylization capability. Besides, our method achieves the best appearance score, showcasing the advanced capability of DiffArtist in capturing high-fidelity appearance details embedded in style prompt. In human evaluations, our method is preferred 67.8% of users, further validating its superiority.
5. Analysis and Discussion
This section provided extended comparisons and discussions. In Sec. 5.1 we provide an in-depth comparison of different style control methods, both qualitative and quantitatively. Sec. 5.2 validate the effectiveness of MLLMs as style evaluator. Sec. 5.3 we provide ablations on delegation and S2A injection. We conclude this section with a discussion in Sec. 5.4.
Figure 5. Comparisons between representative style control methods. DiffArtist offers significantly stronger controllability than other stylization methods. (Left): DiffArtist allows smooth and artistically meaningful structural stylization strength control, while maintaining appearance style strength. (Right): DiffArtist allows fine-grained control of appearance-related style strength while preserving structural and semantic integrity.
5.1. DiffArtist has Fine-grained Controllability
This section aims to validate the unprecedented dual controllability of DiffArtist. Specifically, we provide fine-grained comparisons of the following three groups of representative methods, configured with different control levels: (a) Semantic pyramid: include DiffArtist and DiffArtist implemented with Plug-and-Play structure representation, $f(\mathbf{x})$ (Ours + PnP), and Ctrl-X; (b) Pixel-level map: include InstantStyle (Cui et al., 2023), which is based on ControlNet (Zhang et al., 2023a) to use pixel-level maps as additional conditioning; (c) Noise inversion, which corresponds to the DDIM baseline. For content control in group (a), we set 5 levels, using the following control layers: ($\varnothing$, [1],[1-4],[1-6],[1-8]); for group (b), we evenly sample their respective control strength hyperparameters; and we sample $\tau=[0,5,10,15,20]$ for group (c). Style strength is controlled by sampling CFG weights in [2.5,5,7.5,10] for all groups except for Ctrl-X, which is achieved by adjusting its appearance schedule parameter.
Qualitative Comparison. As visualized in Fig. 5, we observe the strength of structure and appearance style are the most adjustable in DiffArtist, with harmonious, consistent, and disentangled interpolations across a sequence of control levels. Specifically, the compared methods (Ours+PnP) and group (b) primarily produce style at the appearance level, failing to capture the essential fragmented shape of Cubism at varying levels of abstraction. When tuned with different appearance strength, these methods are also incapable of producing artistically meaningful interpolations. While DiffArtist produce gradually bolder strokes, the compared methods only gives more saturated results. This experiment also demonstrate the effectiveness of content representation: the pixel-map based representation can severely violate the semantic integrity, like the example of InstantStyle where the face is barely recognizable. The inversion-based representation, on the other hand, entangles structure and appearance, making it difficult to control one without affecting the other. Further visualization in Appendix. A more clearly demonstrate the strength of DiffArtist.
Quantitative Comparisons. We provide quantitative experiments to further validate the observations from the visualization above. We consider two properties of control: quality and editability. (1) We compare the quality, measured as structure and appearance fidelity scores at different control levels, across different methods. The results are reported in Tab. 2. DiffArtist outperforms others significantly and consistently, demonstrating superior control quality.
We further consider (2) the editability and disentanglement properties of the control within each method. For instance, when increasing structure control, one would expect a gradually stronger structure style, covering a wide range (editability) with minimal influence on the appearance-related style (disentanglement), and vice versa. In other words, during a sequence of control, the structure fidelity should exhibit monotonic growth while the appearance score should resemble a random distribution. To quantify this intuition, we use Spearman’s coefficient of correlation $\rho$ to measure the monotonicity of control, and Kendall’s coefficient of concordance $W$ to quantify the randomness of sequential manipulation. A $\rho$ close to 1 indicates stronger monotonic correlation, while a $W$ close to 0 indicates random ranks. As shown in Tab. 2, the control of DiffArtist is the most editable and disentangled, reaffirming the observation of the visualizations in Fig. 5.
Table 2. Fidelity of structure and appearance control via cross-method comparison. The values corresponds to structure and appearance score (MLLM, $\uparrow$), respectively. Note that the magnitude of scores are only comparable within each column. The best result for each column is in bold.
Method  Structure $\rightarrow$  Appearance $\rightarrow$
lv.1 lv.2 lv.3 lv.4 lv.5 lv.1 lv.2 lv.3 lv.4
Ours 0.62 0.65 0.74 0.63 0.66 0.70 0.74 0.80 0.78
Ours+PnP 0.43 0.38 0.28 0.42 0.34 0.21 0.26 0.27 0.32
Ctrl-X 0.49 0.46 0.36 0.46 0.47 0.41 0.37 0.33 0.42
InstantStyle 0.42 0.42 0.33 0.43 0.35 0.34 0.34 0.35 0.60
DDIM 0.49 0.52 0.52 0.51 0.55 0.59 0.53 0.50 0.52
Table 3. Control editability and disentanglement via inter-method comparison. Higher $\rho$ indicates stronger correlation while lower $W$ means ranks are random. When controlling from one perspective, a high $\rho$ is desired for editability, and a low $W$ for the other aspect is expected for disentangled control. The controls in DiffArtist are the most editable and disentangled.
Sequential Structure-Control Only Sequential Appearance-Control Only
Ours Ours+PnP Ctrl-X InstantStyle DDIM Ours Ours+PnP Ctrl-X InstantStyle DDIM
$\rho(S)\uparrow$ 0.82 0.54 0.32 0.39 0.70 $W(S)\downarrow$ 0.37 0.32 0.36 0.45 0.69
$W(A)\downarrow$ 0.32 0.44 0.45 0.34 0.72 $\rho(A)\uparrow$ 0.71 0.42 0.35 0.26 0.68
5.2. MLLMs are Human-Aligned Style Transfer Metrics
We evaluate how each stylization metric is aligned with human perception by calculating the statistical correlation with rankings from human feedback. To achieve this, we first construct a comparison set of 800 stylized images [^3] The images tested here do not overlap with the experiment in Tab. 1, and compare how human and MLLM preferences correlate.
For human feedback, we consider two groups of users. For the non-expert group, we distributed questionnaires on a crowd-sourced platform, where each participant was required to complete randomly sampled ranking tasks. A total of $n_{1}=200$ participants took part in this study. Filters were applied to validate and clean the raw feedback. We also recruited an expert group of $n_{2}=12$ participants (including the author(s)), who have experience in painting or creative design.
We utilize spearman’s $\rho$ to measure the alignment between metrics and human feedback. The averaged (of all content-style pairs) coefficient $\rho$ for both groups is reported in Tab. 4. As the table shows, the MLLM-based metrics are better aligned with human perception, suggesting its effectiveness as an evaluation metric for style transfer.
Table 4. Metrics correlation with human feedback. We report correlation $\rho$ and combined significance $p$. The VLM-metrics align better with human evaluation for both expert and non-expert users.
Corr. (Non-expert) Corr. (Expert) Metrics $\rho$uparrow p-value $\downarrow$ $\rho$uparrow p-value $\downarrow$ S SSIM (Wang et al., 2004) 0.29 0.12 0.25 0.14 MLLM (GPT-4o) 0.44 0.004 0.34 0.20 MLLM (Gemini v2) 0.42 0.02 0.45 0.03 A CLIP Score 0.05 0.73 0.01 0.75 Pick Score (Kirstain et al., 2023) 0.27 0.11 0.25 0.13 MLLM (GPT-4o) 0.25 0.05 0.22 0.06 MLLM (Gemini v2) 0.48 0.04 0.41 0.02
5.3. Ablations
The delegations enable dual controllability. The control mechanism of DiffArtist is built upon the two delegations. Removing either one will damage the dual controllability. Therefore we compare removing the delegations from full method with the default parameters and summarize the result in Tab. 5. The full methods achieves the best results, demonstrating the synergistic effect of delegations.
Table 5. Ablation on delegation branches. Content delegation reduces unrelated content modification, while the style delegation boosts stylization strength. Controls from both delegations are disentangled from each other.
Metric Method full w/o structure w/o appearance LPIPS $\downarrow$ 0.51 0.76 0.42 CLIP Score $\uparrow$ 25.91 27.69 21.75 Pick Score (Kirstain et al., 2023) $\uparrow$ 20.55 20.57 20.41 Structure (MLLM) $\uparrow$ 0.72 0.37 0.33 Appearance (MLLM) $\uparrow$ 0.62 0.59 0.22
S2A injection promotes uniform spatial distribution of style strength in the style delegation process, thereby avoiding artifacts in the final stylization result. We visualize the denoised style image (from appearance delegation) and the final stylization result in Fig. 6. As observed in the examples, without the S2A injection, the appearance delegation may generate irrelevant structures with uneven appearance distribution, adversely affecting the corresponding regions of the stylization results.
Figure 6. Ablation on S2A injection. S2A injection propagate the high-level semantic to the appearance generation. It avoids spatial-misalignment of appearance-style strength.
5.4. Limitations & Future work
Disentangling the the interplay of structure and appearance in artistic styles is a long-standing challenge. Despite the progress we made in this paper, significant work remains be done. For instance, the structure control in DiffArtist is at a global level, and it cannot control the structure for each object separately. Many art styles exhibit mixed structure variation, like Surrealism and Collage art. We believe that the development of dense structure evaluators with a 2D reward signal is a promising direction (Liang et al., 2024), which would require a deeper understanding of both art and computation.
6. Conclusion
We present the first exploration of structure- and appearance-controllable image stylization. Our contributions include DiffArtist, a generator that fully disentangles structure and appearance during the diffusion process, and a human-aligned evaluator to assess structural and appearance fidelity at the semantic level. Our extensive analysis prove that semantically rich representations are essential for both structure and appearance, allowing for high-fidelity and editability during control, similar to that of a human artist. We hope these design principles broadly inspire future art generation methods to produce truly artistically meaningful paintings.
Appendix A Additional Qualitative Results
Figure 7. Tradeoff between Structure and Appearance Style Control. We present the experimental (solid lines) and simulation (dotted lines) trends of structural and appearance variation in diffusion process. Due to quadratic growth, highly noisy steps are required to achieve strong appearance styles, which are associated with significant structural variation, which can violate semantics. Top: Example stylization results starting from different steps, using the prompt: “Fauvism painting”. When the appearance strength is high ($t=0.8T$), the structure (legs of the horse) are incorrectly modified.
Figure 8. DiffArtist implemented with different diffusion architecture. We implement DiffArtist on the playground-v2 diffusion model. Similar stylization results could be achieved, demonstrating the generalizability of proposed method.
Figure 9. Example of failure cases.
Visualization Additional appearance stylization is in Fig. 10, their source image is in Fig. 11. A grid of different image with different style is in Fig. 12. Additional structure control is in Fig. 13 and Fig. 14.
Additional qualitative comparisons on stylization. We show extended comparison with previous reference-based method in Fig. 15. More qualitative comparisons with existing text-driven image stylization and manipulation methods could be found in Fig. 16 and Fig. 17.
Additional comparisons on fine-grained control. We provide additional comparisons with other control methods in Fig. 18 and Fig. 19. These results clearly demonstrate the advantage of DiffArtist in providing disentangled structural and appearance-level style control. In particular, the Ctrl-X, as an image editing method, produce less visually pleasing result when applied to image stylization. This is because they have a different definition of appearance and structure for the editing of real photos.
Figure 10. Addition results for DiffArtist (with default control parameters). The image semantics are preserved with strong and high-fidelity styles harmoniously integrated.
Figure 11. The corresponding source images of Figure  10.
Figure 12. A grid experiment with different content and styles.
Figure 13. Additional result on structure control - 1.
Figure 14. Additional result on structure control - 2.
Figure 15. Extended comparison with reference-based style transfer methods.
Figure 16. Extended comparison with existing text-driven stylization and manipulation methods.
Figure 17. Extended comparison with existing text-driven stylization and manipulation methods.
Figure 18. Extended comparison on fine-grained structural and appearance-based style control
Figure 19. Extended comparison on fine-grained structural and appearance-based style control
Appendix B On the Structure and Appearance Entanglement in Diffusion Process
B.1. Theoretical analysis
We now explore how the factorization of structure and appearance factorization, defined in Eq. 2, interact and evolve throughout the denoising trajectory $p_{\theta}(\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{T-1},|\mathbf{x}_{T})$. Specifically, suppose the content image $I_{c}$ is inverted into $\mathbf{x}_{1:T}$, inversion-based stylization starts from an intermediate step $\mathbf{x}_{\tau},\tau<T$ for DDIM denoising process. By rearranging Eqn. 1, we obtain:
(7) $\mathbf{x}_{t-1}$ $=\mathcal{A}_{t}\mathbf{x}_{t}+\mathcal{B}_{t}\epsilon_{\theta}(\mathbf{x}_{t},t;y),\quad\mathcal{A}_{t}:=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}$
$\mathcal{B}_{t}$ $:=\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t-1})}}{\sqrt{\alpha_{t}}}$
Based on above formulation, the full stylization process could be expressed as:
(8) $\hat{\mathbf{x}}_{0}=\prod_{j=1}^{T}\mathcal{A}_{j}\cdot\mathbf{x}_{T}$ $+\underbrace{\sum_{k=\tau+1}^{T}\left[\mathcal{B}_{k}\prod_{j=\tau+1}^{k-1}\mathcal{A}_{j}\right]\epsilon^{\prime}(\mathbf{x}_{k-1},k)}_{\text{preserve original structure and appearance}}$
$+\underbrace{\sum_{k=1}^{\tau}\left[\mathcal{B}_{k}\prod_{j=1}^{k-1}\mathcal{A}_{j}\right]\epsilon_{\theta}(\mathbf{x}_{k-1},k;y),}_{\text{generate new structure and appearance}}$
where $\epsilon^{\prime}$ denotes an ideal denoiser that perfectly models the transition distribution $q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$. In Eq. 8, the conceptual denoising term from $\mathbf{x}_{T}$ to $\mathbf{x}_{\tau}$ preserves the structure and appearance in $I_{c}$. The stylization trajectory from $\mathbf{x}_{\tau}$ to $\mathbf{x}_{t}$ is what actually computed, which introduces the desired appearance based on the prompt $y$ but may also lead to uncontrolled structure generation. In this paradigm, the preservation of original appearance and generation of new structure are side-effects that are undesirable. Intuitively, larger $\tau$ leads to a shorter trajectory of $\mathbf{x}_{T}\to\mathbf{x}_{\tau}$, resulting in a stronger appearance in $\hat{\mathbf{x}}_{0}$ with weaker structure preservation (uncontrolled structure stylization), while lower $\tau$ sacrifices stylization strength for stronger structure preservation. In other words, one can not arbitrarily control the strength of appearance and structure without affecting the other.
By combining Eqn. 2 and Eqn. 8, we can quantitatively assess the strength of content preservation and stylization in diffusion process under a particular noise schedule $\alpha_{1:T}$. Specifically, we further assume the SNR of structure and appearance has linear relationship with that of the $\mathbf{x}_{t}$ for $t$:
(9) $\operatorname{SNR}(\mathbf{z}_{t}^{c})\propto\frac{\alpha_{t}}{1-\alpha_{t}};\operatorname{SNR}(\mathbf{z}_{t}^{s})\propto\frac{\alpha_{t}}{1-\alpha_{t}}$
B.2. Simulation
To derive the theoretical trends of structure and appearance strength during the diffusion process, we introduce an additional assumption. Specifically, we assume that the relative significance of each unweighted denoising step $\epsilon(\mathbf{x}_{t},t;y)$ on the final stylized image remains consistent across different timesteps for both structure and appearance. In other words, we assume the SNR of structure and appearance has linear relationship with that of the $\mathbf{x}_{t}$ at time $t$ as characterized by the noise schedule $\alpha_{1:T}$:
(10) $\operatorname{SNR}(\mathbf{z}_{t}^{c})\propto\frac{\alpha_{t}}{1-\alpha_{t}};\operatorname{SNR}(\mathbf{z}_{t}^{s})\propto\frac{\alpha_{t}}{1-\alpha_{t}},$
It is important to note that we do not assume that the relative proportions of structure and appearance are identical at each denoising step.
With above assumption, the effect of varying $\tau$ on structure and appearance of final stylized image could be derived in close-form. In practice, we use the following code to calculate iteratively:
def cum_score(low_t, hi_t, alphas):
res = 0
for k in range(low_t + 1, hi_t):
for j in range(low_t + 1, k - 1):
res += A_t(j, alphas) * B_t(k, alphas)
return res
struct_scores = []
appear_scores = []
for tau in range(0, 50):
crt_struct = cum_score(50 - tau, 50, sampled_alphas)
crt_appear = cum_score(0, tau, sampled_alphas)
struct.append(crt_struct)
appear_scores.append(crt_appear)
B.3. Empirical Result
Due to the inherent inaccuracy of DDIM inversion, the estimation of $\mathbf{x}_{\tau}$ may be imperfect, resulting in unintended modifications in the final sampled image even if no style prompt $y$ is used. To address this issue, we adopt an alternative strategy by randomly sampling 500 Gaussian noise as the $\mathbf{x}_{T}$ of content, which are paired with 500 content prompts. We treat the images that denoised using content prompts for $\tau=T$ steps as the content image, which simulate a perfect inversion technique. To stylize an image, we first denoise the $\mathbf{x}_{T}$ with content prompt for $\tau$ steps, which is subsequently denoised with the style prompt for $T-\tau$ steps. The LPIPS between the stylized image and content image is used as the empirical structural strength, whereas CLIP Deception score (correct classification rate among a set of styles) is used as the empirical appearance strength. The following 10 style prompts is used:
  • ”watercolor style”
  • ”fauvism style”
  • ”pencil sketch style”
  • ”pointillism style”
  • ”art deco style”
  • ”impressionism style”
  • ”surrealism style”
  • ”pop art style”
  • ”cubism style”
  • ”abstract expressionism style”
The result for both simulation and empirical results are in Fig. 7. The result show a good fit, and it turns out that the structure modification appears to be linear, with the stylization strength is quadratic with respect to $\tau$. Moreover, this result further evidenced the issue of S-A entanglement in the diffusion process.
Appendix C Details on MLLM-based metrics
C.1. Implementation Details
The stylized images, style prompt, and the instruction prompts are feed to MLLM for inference. We compose stylized images as a grid image with numbers at the top-left corner. The full prompt template for structure and appearance score is available in Tab. 6.
C.2. Correlation with Human Preference
Human Question Collection We distributed the questionnaire on a crowd-sourcing platform, where each participant was required to complete up to 20 randomly sampled ranking tasks. An example of the user interface is provided in Fig. 20. A total of 200 participants took part in this study. To ensure the validity of the responses, we included attention-check questions. If a participant answers an attention-check question incorrectly, all of their responses will be marked invalid. Response that are made with less than 20 seconds are also removed.
Table 6. Prompt templates for MLLM-based metrics. [IMG], [STYLE] and [NUM_METHOD] is the placeholder for combined image, style and number of methods, respectively.
Structure Score Appearance Score
”[IMAGE] A content (source) image (top left) and [NUM_METHOD] stylized images in the style of [STYLE] are placed as a grid. ” ”The stylized images are indexed from left to right, and from top to bottom. ” ”Compare, analyze and discriminately rank the fidelity to which the structure described in the style of [STYLE] is transferred to the source image.” ”You should focus on the fidelity of structure-related style component only, such as the lines, shapes, geometry, layout, and perspective. You should not consider the style related to appearance (e.g., texture, color, stroke, and pattern). You should also consider how the structure of [STYLE] is integrated with the source image.” ”Stylized image that has (1) limited style strength, (2) structure that is mis-aligned with the style, or (3) significant artifacts and distortions of the semantic integrity (e.g., the original object and scene become unrecognizable) unless the distortion is explicitly intended by the style of [STYLE], and (4) un-harmonious integration with the source image should be considered of in lower rank. In other words, if a stylized image is not an artistically meaningful painting of the source image in target style, then it should be rated lower. Images that harmoniously integrate the structure of [STYLE] with the source image should be rated higher.” ”Rank the [NUM_METHOD] images in ascending order from 1 to [NUM_METHOD], where the highest rank of [NUM_METHOD] means the best structural fidelity. No images shall have the same ranking. ” ”As an expert in art, return your thinking in short (what structure is desired, and how the ranking is decided for each image in short), and ranks for each image id in a Python Dict, [’thinking’:str, ’rank’:List[[NUM_METHOD]]]. Do not include any other string in your response.” ”[IMAGE] A content (source) image (top left) and [NUM_METHOD] stylized images in the style of [STYLE] are placed as a grid. ” ”The stylized images are indexed from left to right, and from top to bottom. ” ”Compare, analyze and discriminately rank the fidelity to which the appearance described in the style of [STYLE] is transferred and to the source image.” ”You should focus on the fidelity of appearance-related art style component only, such as the texture, color, stroke, and pattern. Note that it does not simply means color pallette and saturation. You should not consider the style related to structure (e.g., lines, shapes, geometry, layout, and perspective), unless the original scene become unrecognizable. You should also consider how the appearance of [STYLE] is integrated with the source image.” ”Stylized image that has (1) limited style strength, (2) visual appearance that is mis-aligned with the style, (3) significant artifacts and distortions of the semantic integrity (e.g., the original object and scene become unrecognizable) unless the distortion is explicitly intended by the style of [STYLE] and (4) un-harmonious integration with the source image should be considered of in lower rank. In other words, if a stylized image is not an artistically meaningful painting of the source image in target style, then it should be rated lower. Images that harmoniously integrate texture, color, stroke, and pattern should be rated higher.” ”Rank the [NUM_METHOD] images in ascending order from 1 to [NUM_METHOD], where the highest rank of [NUM_METHOD] means the best appearance fidelity. No images shall have the same ranking.” ”As an expert in art, return your thinking (what appearance is desired, and how the ranking is decided for each image in short), and ranks for each image id in a Python Dict, [’thinking’:str, ’rank’:List[[NUM_METHOD]]]. Do not include any other string in your response.”
Figure 20. Example user interface in collecting human preference. The system will prevent user from selecting the same ranking.
Appendix D Additional Discussion and Analysis
D.1. Generalizability of DiffArtist
To demonstrate the generalizability of DiffArtist across different U-Net-based diffusion architectures, we implement our method on Playground v2444https://huggingface.co/playgroundai/playground-v2-1024px-aesthetic, which utilizes the SDXL architecture, distinct from Stable Diffusion 2.1. Several results are provided in Fig. 8. These results validate that DiffArtist serves as a versatile control method applicable to various U-Net-based diffusion models, regardless of their underlying architectural differences.
D.2. Additional Ablations on S2A Injection
Figure 21. Ablation Study on S2A Layers $S_{S2A}$. Increasing the number of S2A layers compels the appearance delegation to generate higher-frequency style features (stokes, points) while diminishing low-frequency tonal components (large color fields). Empirically, our default configuration $[1,2]$ strikes an optimal balance between enhancing style detail and preserving essential content structure.
In this section we provide additional ablations to study the effect of proposed S2A design.
Ablation on S2A layers $S_{S2A}$. We ablate the number of injection layer used in the S2A injection (i.e., $S_{S2A}$). As illustrated in Fig. 21, the S2A layers influence the frequency bands of style delegation. Incorporating only early layers (e.g., [1, 2]) focuses on generating low-frequency style features such as tones and small objects, while adding more layers facilitates the creation of high-frequency style details like stroke shapes. Empirically, we set the S2A injection layers to [1, 2] by default, as using additional layers typically results in blurriness in the stylized outputs.
D.3. Failure Case
In our experiment, we identify a rare ($<$1%) and special failure case in the proposed methods. Specifically, for certain content image, its stylization result will consistently contains black and white chessboard-pattern artifacts. We provide one example in Fig. 9.

### v4

@@
 Abstract.
-Artistic style includes both structural and appearance elements.
-Existing neural stylization techniques primarily focus on transferring appearance features such as color and texture, often neglecting the equally crucial aspect of structural stylization.
-In this paper, we present a comprehensive study on the simultaneous stylization of structure and appearance of 2D images.
-Specifically, we introduce DiffArtist, which, to the best of our knowledge, is the first stylization method to allow for dual controllability over structure and appearance.
-Our key insight is to represent structure and appearance as separate diffusion processes to achieve complete disentanglement without requiring any training, thereby endowing users with unprecedented controllability for both components.
-The evaluation of stylization of both appearance and structure, however, remains challenging as it necessitates semantic understanding.
-To this end, we further propose a Multimodal LLM-based style evaluator, which better aligns with human preferences than metrics lacking semantic understanding.
-With this powerful evaluator, we conduct extensive analysis, demonstrating that DiffArtist achieves superior style fidelity, editability, and structure-appearance disentanglement.
-These merits make DiffArtist a highly versatile solution for creative applications.
-Project homepage: https://github.com/songrise/Artist.
-Generative art, Text-driven stylization, Structure and appearance, Stylization evaluation, Multimodal LLM applications,
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Arxiv; 2025; NYisbn: 978-1-4503-XXXX-X/2018/06submissionid: ccs: Computing methodologies Image processingccs: Applied computing Fine artsccs: Computing methodologies Artificial intelligence
+Artistic styles are defined by both their structural and appearance elements.
+Existing neural stylization techniques primarily focus on transferring appearance-level features such as color and texture, often neglecting the equally crucial aspect of structural stylization.
+To address this gap, we introduce DiffArtist, the first 2D stylization method to offer fine-grained, simultaneous control over both structure and appearance style strength.
+This dual controllability is achieved by representing structure and appearance generation as separate diffusion processes, necessitating no further tuning or additional adapters.
+To properly evaluate this new capability of dual stylization, we further propose a Multimodal LLM-based stylization evaluator that aligns significantly better with human preferences than existing metrics.
+Extensive analysis shows that DiffArtist achieves superior style fidelity and dual-controllability compared to state-of-the-art methods.
+Its text-driven, training-free design and unprecedented dual controllability make it a powerful and interactive tool for various creative applications.
+Project homepage: https://diffusionartist.github.io.
+Generative art; Text-driven stylization; Structure and appearance; Stylization evaluation; Multimodal LLM applications
+copyright: acmlicensedjournalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755010isbn: 979-8-4007-2035-2/2025/10submissionid: $c$s: Computing methodologies Appearance and texture representationsccs: Computing methodologies Image manipulationccs: Applied computing Fine arts
 1. Introduction
-Figure 1. DiffArtist enables disentangled and fine-grained control of style strength from two orthogonal perspectives: structure and appearance. The prompt is “The Dream, by Picasso.”
-The essence of an artistic style lies not only in its visual appearance but also the structural elements (Gombrich and Gombrich, 1995; Jung et al., 2024).
+Figure 1. DiffArtist enables disentangled and fine-grained control of style strength from two orthogonal perspectives: structure and appearance. The style prompt is “The Dream, by Picasso.”
+The essence of an artistic style lies not only in its appearance—color and texture—but also its structure—geometry and composition (Gombrich and Gombrich, 1995; Jung et al., 2024; Goodman, 1976).
 For example, the fragmentation of figures in Picasso’s Cubist works and the undulating sky in Van Gogh’s “Starry Night”, each contributing distinctly to their artistic expression.
-To automate the generation of artworks, existing neural stylization approaches (Huang and Belongie, 2017; Deng et al., 2022; Kwon and Ye, 2022; Wang et al., 2022; Huang et al., 2024; Cui et al., 2023) have predominantly focused on manipulating appearance-level attributes such as color and textures.
-The structural elements in the source image, however, are often modeled as the “content” and are explicitly preserved (Wang et al., 2024b; Wen et al., 2023; Yang et al., 2022).
-Consequently, these methods typically lack the capacity to curate meaningful structural-styles, substantially limiting their artistic potential and customizability.
-Unlike appearance, structural stylization is inherently much more challenging as it demands artistically meaningful variations that (1) are aligned with the target style, (2) harmoniously integrated with the original image, and (3) without violating the semantic integrity of the source content.
-These objectives are at a higher, potentially semantics-level than the appearance-style, and, therefore, are less readily achieved.
-While descriptors like AdaIN (Huang and Belongie, 2017) and Gram loss (Gatys et al., 2015) may suffice for appearance-style modeling, the lack of adequate structure representation and a structure-style evaluator represent major obstacles in the development of structural stylization techniques.
-This challenge becomes even more pronounced in contemporary multimodal generation scenarios, where the style as textual prompt lacks a direct visual counterpart (Huang et al., 2024; Wang et al., 2023a; Cui et al., 2023; Hertz et al., 2024).
-In search of the dual controllability of structure and appearance elements, the recent development of image generation models, in particular the diffusion models (DMs) has provided a promising solution.
-Compared with traditional stylization, the generative process allows for the sampling of stylized images of greater diversity in both structure and appearance.
-Building upon DMs, the stylization problem can re-framed as controlled image generation, where the reference image and style are additional conditions for sampling of stylized images.
-However, achieving this control in an artistically meaningful way is far from straightforward, as the diffusion process inherently entangles the generation of structure and appearance, as analyzed in this paper.
-This entanglement presents a major challenge for full controllability, resulting in a tradeoff between structural and appearance style strength (i.e., the S-A tradeoff).
-Consequently, existing diffusion-based approaches are either prone to undesired content modification (Chung et al., 2024; Lyu et al., 2023; Wang et al., 2024b) or suffer from constrained appearance-style strength (Wang et al., 2023b).
+Existing neural stylization approaches (Huang and Belongie, 2017; Deng et al., 2022; Kwon and Ye, 2022; Wang et al., 2022; Huang et al., 2024; Cui et al., 2023) predominantly focus on manipulating appearance-level attributes.
+The structural elements in the source image, however, are often viewed as part of “content” and are explicitly preserved (Wang et al., 2024b; Wen et al., 2023; Yang et al., 2022, 2023).
+This fundamental limitation prevents them from capturing the true essence of an art style, severely restricting their expressive potential and customizability.
+The root of this limitation lies in the inherent complexity of structural stylization.
+Unlike appearance-style transfer, structural stylization requires a delicate balance between three competing objectives: (1) aligning with the target style, (2) harmonizing with the source image’s composition, and (3) preserving the core semantic integrity of the content.
+These objectives operate at a high-level semantic plane, exposing a critical gap in current methods.
+While descriptors like AdaIN (Huang and Belongie, 2017) and Gram loss (Gatys et al., 2015) may suffice for appearance-style modeling, the lack of adequate structure representation and structure-style evaluators presents significant obstacles in the development of structural stylization techniques.
+This challenge is amplified in recent multimodal generation scenarios, where a style prompt offers no explicit visual template (Huang et al., 2024; Wang et al., 2023a; Cui et al., 2023; Hertz et al., 2024).
+The advent of Diffusion Models (DMs) offers a powerful new paradigm for achieving this dual controllability, as their generative sampling process enables far greater structural and appearance diversity than prior methods.
+This reframes stylization as a conditional generation task, guided by a source image and a style prompt (image or text).
+However, this generative power comes with a critical, unaddressed challenge: the diffusion process inherently entangles the generation of structure and appearance.
+We identify this as a fundamental Structure-Appearance (S-A) Tradeoff: intensifying structural changes inadvertently corrupts appearance style, while strengthening appearance washes out structural transformations.
+This tradeoff directly explains the core failures of existing diffusion-based methods, which are either prone to severe content degradation (Chung et al., 2024; Lyu et al., 2023; Wang et al., 2024b) or suffer from weak, constrained stylization (Wang et al., 2023b; Yang et al., 2023).
 Achieving dual controllability in the stylization thus remains an open question.
-In this paper, we present the first comprehensive study for disentangled structure and appearance control for 2D stylization.
-We first present an styler, DiffArtist, which, to the best of our knowledge, is the first to enable the dual controllability for text-driven stylization.
-DiffArtist is based on explicit disentanglement of the structural and appearance generation in the diffusion process.
-It is plug-in with U-Net-based DMs, requiring neither optimization nor additional adapters like ControlNet (Zhang et al., 2023a) or IP-Adapter (Ye et al., 2023) (i.e., zero-shot), making it highly efficient for user personalization.
-This superior controllability overcomes the S-A tradeoff in DMs.
-Compared with conventional methods and other DM-based methods (e.g, (Wang et al., 2024a; Lin et al., 2024; Zhang et al., 2023a)), we demonstrate DiffArtist achieves significantly better dual controllability without deterioration in semantics, as exemplified in Fig.
-1.
-Aside from structural stylization, it also allows appearance stylization with un-represented high fidelity, when configured with strict structure preservation, thereby unleashing full customization power.
-In addition to the proposed styler, we further explore the development of art evaluators with human-aligned perceptions of structure and appearance style.
-The challenges for this exploration are threefold: (1) Structural style components are at a higher level than appearance and cannot be evaluated in isolation without considering their integration with content.
-(2) Evaluators must possess contextual awareness to detect structural artifacts that compromise semantic coherence.
-(3) Evaluators should also effectively associate textual style instructions with corresponding visual features for cross-modal evaluation.
-Multimodal LLMs (MLLMs), known for their robust zero-shot semantic-level understanding capabilities, emerge as an optimal solution to address these challenges.
-To this end, we propose (to the best of our knowledge) the first MLLM-based metrics for evaluating the structure and appearance in artistic stylization.
-We empirically demonstrate that this metric aligns significantly better with human artistic perception than existing metrics (Kirstain et al., 2023; Wang et al., 2004; Radford et al., 2021), paving the way for more human-aligned style evaluation.
+To solve this, we introduce DiffArtist, the first framework to our knowledge that offers explicit, disentangled control over both structure and appearance in 2D stylization.
+At its core, DiffArtist explicitly disentangles the structural and appearance generation as separated diffusion processes, with shared semantic information.
+This design directly overcomes the fundamental S-A tradeoff and functions as a zero-shot, plug-and-play module for any pretrained U-Net-based DM, requiring no costly fine-tuning or external adapters (Zhang et al., 2023a; Ye et al., 2023).
+As evidenced in this paper, this design provides true disentanglement—a key advantage over ControlNet-based methods (Wang et al., 2024a; Lin et al., 2024) where adjusting one style factor adversely impacts the other.
+As demonstrated in Fig.
+1, this unprecedented level of control allows DiffArtist to achieve strong, semantically coherent stylization, unleashing the full creative potential of dual-style customization.
+Evaluating this novel capability of dual control requires understanding on image semantics, where existing evaluation metrics obsolete.
+This exposes a critical need for a new evaluation paradigm.
+To address this, we introduce our second major contribution: a Multimodal LLM (MLLM)-based evaluator designed for dual stylization.
+We argue that any such evaluator must satisfy three key criteria: (1) operate at a high-level semantic plane to assess structure, (2) possess contextual awareness to maintain semantic integrity, and (3) perform robust cross-modal association between text prompts and visual forms.
+By leveraging the zero-shot reasoning of MLLMs, our proposed metric meets these criteria.
+We empirically show that it aligns significantly better with human artistic judgment than existing stylization metrics (Kirstain et al., 2023; Wang et al., 2004; Radford et al., 2021), establishing a more reliable and human-centric standard for future stylization research.
 We summarize our contributions as follows:
   1. (1)
-    We propose DiffArtist, the first stylization method that enables the dual controllability of structure and appearance-related style.
+    We identify the S-A tradeoff in diffusion models as the key challenge for disentangled dual controllability.
   2. (2)
-    We present a novel MLLM-based evaluator for evaluation of structure and appearance in artistic stylization, which aligns better with human perception.
+    We propose DiffArtist, the first 2D stylization method that enables the dual controllability of structure and appearance.
   3. (3)
-    Extensive experiments demonstrate DiffArtist achieves superior stylization fidelity, with highly editable and disentangled control.
+    We present a novel MLLM-based evaluator for evaluating structure and appearance in artistic stylization, which aligns better with human perception.
+  4. (4)
+    Extensive experiments demonstrate that DiffArtist achieves superior stylization fidelity, control editability, and disentanglement than existing approaches.
 2. Related Works
+Figure 2.
+Overview of DiffArtist.Our method disentangles stylization by processing structure and appearance through two independent diffusion trajectories (delegations).
+At each denoising step, the main stylization branch is conditioned on semantic-level features from the structure and appearance delegations.
+All three branches share the same pretrained U-Net parameters, and perform full denoising of $\tau$ steps.
+The entire framework operates without requiring any fine-tuning or adapters.
 The Stylization of Structure and/or Appearance.
 Structure and appearance collectively define the style of a visual representation.
-Existing neural stylization methods (Huang and Belongie, 2017; Deng et al., 2022; Kwon and Ye, 2022; Wang et al., 2022; Huang et al., 2024; Cui et al., 2023) has predominately focused on appearance stylization, generally achieved via encoder-decoder architecture.
+Existing neural stylization methods (Huang and Belongie, 2017; Deng et al., 2022; Kwon and Ye, 2022; Wang et al., 2022; Huang et al., 2024; Cui et al., 2023) have predominantly focused on appearance stylization, generally achieved via an encoder-decoder architecture.
 Only a few papers (Kim et al., 2020; Zhou et al., 2024) focus on transferring structural style components between 2D images.
-This can be achieved via image deformation.
-For instance, one may calculate a correspondence between the structure of content and style image, to perform a non-rigid deformation.
-However, these methods usually operate images in specific domains like portrait, requiring a stylized image as reference.
-They also do not allow for dual controllability.
-The dual controllability has recently been explored in image editing literature (Lin et al., 2024), yet they focus on targeted editing of realistic photos at single control level, instead of harmonious transfer between photo and artworks with high editablity.
-The structure stylization is more widely explored in 3D stylization, which has an explicit formulation of shapes (e.g, vertices (Michel et al., 2022), density (Jung et al., 2024)) and appearance (e.g., texture (Chen et al., 2023b) map, radiance (Wang et al., 2023a)), yet they cannot be directly applied to the 2D case.
+This is usually accomplished by calculating a correspondence between content and style image and performing a non-rigid deformation.
+However, these methods typically operate on images in specific domains like portraits, requiring an in-domain stylized reference.
+Moreover, they do not enable dual controllability.
+Very recently, dual controllability has been explored in Ctrl-X (Lin et al., 2024), but its framework is designed for localized, image edits rather than the global, harmonious style transformations.
+Meanwhile, the concept of disentangled control is explored in 3D synthesis, where explicit representations of shape (e.g., meshes (Michel et al., 2022) and radiance fields (Wang et al., 2023a)) and appearance (e.g., texture mapping (Chen et al., 2023b)) make separation natural.
+The success in 3D, however, is not directly transferable to the 2D domain due to the lack of explicit geometric priors in single images.
 In this paper, we explore the first dual controllable 2D image stylization method.
@@
 Text-driven image stylization aims to stylize a source image according to style prompts.
-Early methods achieve it by optimizing certain image representation (Michel et al., 2022; Kwon and Ye, 2022; Wang et al., 2023a; Gal et al., 2022; Kim et al., 2022) with a multimodal similarity loss, usually implemented as the CLIP loss (Radford et al., 2021).
+Early methods achieve it by optimizing certain image representation (Michel et al., 2022; Kwon and Ye, 2022; Wang et al., 2023a; Gal et al., 2022; Kim et al., 2022) with a multimodal alignment objective, typically implemented as the CLIP loss (Radford et al., 2021).
 Recently, it was discovered that text-to-image (T2I) DMs could also be adapted for similar optimization schemes (Poole et al., 2022; Hertz et al., 2023; Jiang et al., 2023; Kawar et al., 2023).
-These optimization-based methods are costly and slow, motivating recent methods explore the feed-forward paradigm.
-Instruct-Pix2Pix (Brooks et al., 2023) tune the diffusion model with a language model for generalized editing tasks.
+These optimization-based methods are costly and slow, motivating recent exploration of the feed-forward paradigm.
+Instruct-Pix2Pix (Brooks et al., 2023) tunes the diffusion model along with a language model for generalized editing tasks.
 Diffstyler (Huang et al., 2024) learn a content and style-specific denoiser for disentanglement.
-FreeStyle (He et al., 2024) modulate U-Net feature for training-free stylization.
-Concurrent with this exploration is the stylized image generation.
-While related, they focus a different setting where the style is extracted from an image and the content is a prompt (Hertz et al., 2024; Wang et al., 2024a, b; Gao et al., 2024; Chen et al., 2023a).
+FreeStyle (He et al., 2024) modulates the U-Net feature for training-free stylization.
+Concurrent with this exploration is the stylized image generation (Hertz et al., 2024; Wang et al., 2024a, b; Gao et al., 2024; Chen et al., 2023a).
+While related, they focus on a different setting where the style is extracted from an image and the content is a prompt.
 In this work, we focus on the structure and appearance control in text-driven stylization scenarios.
 Quantitative Evaluation of Style Transfer.
-Traditional metrics such as Gram Loss (Gatys et al., 2016), LPIPS (Zhang et al., 2018), and Fréchet Inception Distance (FID) (Heusel et al., 2017) evaluate quality of stylization from various perspectives.
-These metrics are at a relatively low-level, and are found to be misaligned with the holistic human artistic perception (Bo et al., 2018; So, 2023; Ioannou and Maddock, 2024; Chen et al., 2024; Yi et al., 2023; Wright and Ommer, 2022).
-To build art-specific evaluators, recent methods like ArtFID (Wright and Ommer, 2022) and ArtScore (Chen et al., 2024) attempt to quantify ”artness” of images, yet these methods cannot handle the open-vocabulary text-driven stylization, and they do not evaluate structure and appearance separately.
-In this paper, we propose a semantic-level MLLM-based evaluator to assess the structure and appearance fidelity, which aligns better with human perception.
+Quantitatively evaluating style transfer is a long-standing problem.
+Initial approaches repurposed low-level metrics, including Gram Loss (Gatys et al., 2016), LPIPS (Zhang et al., 2018), and FID (Heusel et al., 2017).
+However, recent literature found that these metrics are fundamentally incapable of capturing the holistic and semantic qualities of human artistic perception (Bo et al., 2018; So, 2023; Ioannou and Maddock, 2024; Chen et al., 2024; Yi et al., 2023; Wright and Ommer, 2022).
+In response to these shortcomings, art-specific evaluators like ArtFID (Wright and Ommer, 2022) and ArtScore (Chen et al., 2024) were developed to better quantify the abstract concept of ”artness”.
+Nevertheless, they cannot handle open-vocabulary text-driven stylization and lack the mechanism to evaluate structure and appearance separately.
+This paper propose a semantic-level MLLM-based evaluator to assess the structure and appearance fidelity, which aligns better with human perception.
 3. Methodology
-Figure 2.
-Overview of DiffArtist.
-We disentangle the structure and appearance of image as separate diffusion processes.
-All branches share the same U-Net parameters, and perform full denoising of $\tau$ steps.
-The feature pyramid from structure delegation is injected into the main branch for structure control, and the semantic-aware appearance is controlled via AdaIN.
-3.1. Objective: Dual Controllability
-Given a content image $I$, a style prompt $y$, and a U-Net-based DM $\mathcal{G}$, our objective is to control $\mathcal{G}$ to derive a distribution of stylized images $\hat{I}\sim\mathcal{G}(I,y)$ that harmoniously compose the provided content and style.
-Specifically, we aim to achieve disentangled controllability in both structure and appearance-related style.
-Our definition of structure and appearance in 2D imagery is mainly based on fine-art (Gombrich and Gombrich, 1995).
-A formal definition of them relates to visual semiotics (Van Leeuwen and Jewitt, 2000) and extend beyond the scope of this paper.
-Generally speaking, structure corresponds to the geometry, like contours and curvatures, while appearance corresponds to patterns like strokes and color pallette.
+3.1. Objective: Disentangled Dual Controllability
+Given a source image $I$, a text-based style prompt $y$, and a pretrained diffusion model $\mathcal{G}(\cdot)$, our primary objective is to generate a stylized image $\hat{I}$ that preserves the semantic content of $I$ while harmoniously embodying the style described by $y$.
+The core innovation we pursue is disentangled dual control, meaning that decomposing the style prompt $y$ into two orthogonal components—structure and appearance—and controlling their strength independently.
+Our definition of structure and appearance in a 2D image is mainly based on fine art (Gombrich and Gombrich, 1995).
+A formal definition of them is challenging as it relates to visual semiotics (Van Leeuwen and Jewitt, 2000; Goodman, 1976), extending beyond the scope of this paper.
+Generally speaking, structure corresponds to the shapes, like contours and curvatures, while appearance corresponds to local patterns, like strokes and color palettes.
 We also aim to develop an evaluator $\mathcal{E}$, which can evaluate the fidelity of structure and appearance style in a way aligned with human perception.
-The remaining parts of this section are organized as below.
+The remaining parts of this section are organized as follows.
 In Sec.
@@
 In Sec.
-3.33.4 we explain the motivation and design of control at a high-level, and details are described in Sec.
+3.33.4, we explain the motivation and design of control at a high level, and details are described in Sec.
 3.5.
@@
 3.2. Preliminary: DDIM Inversion
-To stylize a real image $\mathbf{x}_{0}:=I$ using DMs, inversion-based methods first approximate the noise latent $\mathbf{\tilde{x}}_{1:T}$ of it, typically using techniques such as DDIM inversion (Song et al., 2020).
-With the inverted latents, stylization can be achieved through generation with altered conditions (usually given as a style prompt $y$).
+To stylize a source image $\mathbf{x}_{0}:=I$ using DMs, inversion-based methods first approximate the noise latents $\mathbf{x}_{1:T}$ of $I$, achieved via techniques such as DDIM inversion (Song et al., 2020).
+Stylization is then performed through re-generation with altered conditions (usually specified as a style prompt $y$).
 Specifically, one may start with an intermediate step $\mathbf{x}_{\tau}$ (i.e., control point), where $\tau\in[1,T]$ for iterative DDIM sampling.
-Each denoising step is defined as:
+Each denoising step is formulated as follows:
 $\mathbf{x}_{t-1}$ $=\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t;y)}{\sqrt{\alpha_{t}}}\right)$
 (1) $\qquad+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(\mathbf{x}_{t},t;y),$
-where $t$ is the timestep, and $\alpha_{1:T}$ is a specific noise schedule.
-The assumption in this paradigm is that with a proper $\tau$, the resulting stylized image $\hat{I}:=\hat{\mathbf{x}}_{0}$ harmoniously integrate the content of original image ${\mathbf{x}_{0}}$ with the style in prompt $y$.
+where $\epsilon_{\theta}$ is the denoiser, $t$ is the timestep, and $\alpha_{1:T}$ is a predefined noise schedule.
+The assumption of this paradigm is that with a proper $\tau$, the resulting stylized image $\hat{I}:=\hat{\mathbf{x}}_{0}$ harmoniously integrates the structure and appearance of style in prompt $y$ with the source image $I$.
 3.3. Structure and Appearance in Noise Space
-Existing neural stylization methods generally view an image as a composition of certain “content” and “style” components  (Li et al., 2017; Huang and Belongie, 2017; Wang et al., 2023b; Kotovenko et al., 2019; Sanakoyeu et al., 2018; Zhang et al., 2023b, 2020), where the style usually refers to the low-level visual features such as appearance.
-As this paper motivates to stylize both structure and appearance, we adopt a different modeling to decompose image as appearance and structure components: $\mathbf{x}_{0}=\mathcal{G}_{0}(\mathbf{z}^{a}_{0},\mathbf{z}^{s}_{0})$, where $\mathcal{G}_{0}=(\cdot,\cdot)$ is a composition function, $\mathbf{z}^{a}_{0}$ and $\mathbf{z}^{s}_{0}$ are the latent appearance and structure factorization, respectively.
-This is a “static” perspective.
-In the diffusion process, the distribution of image $x_{0}$ is tied with the intermediate distributions in $\mathbf{x}_{1:T}$, where a denoiser $\epsilon_{\theta}$ learns the transition $q_{\theta}(x_{t-1}|x_{t},y)$ via $\epsilon$-prediction.
-We therefore posit similar factorization of noise-residuals $\epsilon_{\theta}$:
-(2) $\epsilon_{\theta}(\mathbf{x}_{t},t;y)=\mathcal{G}_{t}\left(\kappa_{t},\psi_{t}\right),$
-where $\kappa_{t}$ and $\psi_{t}$ denote the structure and appearance representation at diffusion time-step $t$ (detailed later), and $\mathcal{G}_{t}(\cdot,\cdot)$ is the noise-space composition function at time $t$.
-3.4. Structure and Appearance as Delegate Process
-Controlling the structure and appearance-level stylization in DMs presents unique challenges due to inherent entanglement.
-We provide both theoretical and empirical evidence for this phenomenon in Appendix.
-B, revealing the technical challenge of S-A tradeoff in DM.
-The key insight from this analysis is that solidary control point $\mathbf{x}_{\tau}$ used in existing DMs-based stylization does not suffice for S-A disentanglement.
-Consequently, we argue that the structure and appearance shall be represented and controlled over complete and separate diffusion trajectories.
-We leverage two additional diffusion processes with shared information for dual controllability, which we call delegate branches.
-We initialize the structure and main branch from the inverted noise, while appearance delegation starts from a sampled noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
-With these delegations, we are able to control the stylization over the entire diffusion process.
-The controlled main stylization branch can be denoted as:
-(3) $\epsilon^{m}_{\theta}(\mathbf{x}_{t},t;y,\kappa^{s},\psi^{a})=\mathcal{G}_{t}\left(\kappa^{s}\circ\kappa^{m},\psi^{a}\diamond\psi^{m}\right),\quad t\in[1,T]$
-where the superscripts $s$, $a$, and $m$ denote the factorization extracted from the structure, appearance delegation, or main branch, respectively.
+Prevailing neural stylization methods are built on a paradigm that separates an image into “content” and “style” (Li et al., 2017; Huang and Belongie, 2017; Wang et al., 2023b; Kotovenko et al., 2019; Sanakoyeu et al., 2018; Zhang et al., 2023b, 2020; Ding et al., 2024).
+In this view, the style usually refers to the feature maps extracted from certain layers of a neural network.
+To advance stylization towards both structure and appearance controllability, we adopt different modeling that decomposes an image as its structure and appearance components: $\mathbf{x}_{0}=\mathcal{G}_{0}(\mathbf{z}^{s}_{0},\mathbf{z}^{a}_{0})$, where $\mathcal{G}_{0}=(\cdot,\cdot)$ is a composition function, $\mathbf{z}^{s}_{0}$ and $\mathbf{z}^{a}_{0}$ are the latent structure and appearance factorization, respectively.
+This is a “static”, image-level perspective.
+In the diffusion process, the distribution of image $x_{0}$ is tied with the intermediate distributions in $\mathbf{x}_{1:T}$, where the denoiser $\epsilon_{\theta}$ learns the transition $q_{\theta}(x_{t-1}|x_{t},y)$ via $\epsilon$-prediction.
+Therefore, we posit similar factorization of predicted noise residual, which is a “dynamic” decomposition across the full frequency bands:
+(2) $\epsilon_{\theta}(\mathbf{x}_{t},t;y)=\mathcal{G}_{t}\left(\kappa_{t},\psi_{t}\right),\quad t\in[0,T]$
+where $\kappa_{t}$ and $\psi_{t}$ denote the structure and appearance representation at diffusion time-step $t$ (detailed later), respectively.
+$\mathcal{G}_{t}(\cdot,\cdot)$ is a conceptual noise-space composition function at time $t$.
+3.4. Structure and Appearance as Delegate Diffusion Process
+We argue that the fundamental obstacle to achieving dual control in diffusion-based stylization is the inherent entanglement of structure and appearance.
+Our analysis, detailed in Appendix D, pinpoints the source of this problem: the reliance on a single latent trajectory $\mathbf{x}_{\tau}\xrightarrow{}\mathbf{x}_{0}$.
+This monolithic generation process forces structural and appearance attributes to compete for influence at every denoising step, creating the S-A tradeoff that fundamentally limits controllability.
+To break this bottleneck, we propose a novel mechanism that stylizes an image with separate diffusion trajectories, as illustrated in Fig.
+2.
+Specifically, we leverage two supplementary diffusion processes with shared information, called delegate branches.
+We initialize the structure and main branch from the inverted noise $\mathbf{x}_{T}$, while appearance delegation starts from a Gaussian $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
+These delegations enable controlling the stylization over the entire diffusion process.
+The controlled main branch can be denoted as:
+(3) $\epsilon^{m}_{\theta}(\mathbf{x}_{t},t;y,\kappa^{s}_{t},\psi^{a}_{t})=\mathcal{G}_{t}\left(\kappa^{s}_{t}\circ\kappa^{m}_{t},\psi^{a}_{t}\diamond\psi^{m}_{t}\right),\quad t\in[0,T]$
+where the superscripts $s$, $a$, and $m$ denote the factorization extracted from the structure, appearance delegation, and main branch, respectively.
 The $\circ$ and $\diamond$ are two non-commutative control operators.
 3.5. Structure and Appearance Representations in Denoising U-Net
-In this section we formulate the $\kappa$ and $\psi$ in U-Net-based denoiser for enabling structure and appearance disentangled style control.
-Pyramidal Structure Representation.
-A multi-scale, and semantic-rich image representation is essential for structure stylization at different level of abstractness.
-To achieve this, we find the hidden features in the ResBlock of denoising U-Net encode appearance-invariant image semantics across varying $t$ (see Fig.
+Having established the control mechanism in Eq. 3, we now formulate the $\kappa$ and $\psi$ in a U-Net-based denoiser for disentangling structure and appearance control.
+Pyramidal Structure Representation $\kappa$.
+To effectively control structural stylization, we require a representation that captures image semantics at multiple levels of abstraction.
+We identify the hidden features in the ResBlock of denoising U-Net as the ideal substrate for this purpose, which robustly encode appearance-invariant image semantics across varying $t$ (see Fig.
 3).
-Formally, we denote the hidden feature of ResBlock as $h_{i}(\mathbf{x}_{t})$, where $i\in\{1,2,\ldots,N_{res}\}$ indexes the ResBlocks in the decoder with varying spatial resolution up to $N_{res}$.
-Stacking the hidden feature from all layers forms a feature pyramid as the structure representation of $\mathbf{x}_{0}$ at $t$:
-(4) $\kappa_{t}$ :=$:=\{h_{i}(\mathbf{x_{t}})\}_{i\in S_{res}},$
-[…]
-and $\text{and }S_{res}$$\subseteq\{1,2,\ldots,N_{res}\}.$
-Technically, the representation above captures the semantics of $x_{0}$ at multiple spatial resolutions and (almost) full signal-to-noise ratio (SNR), making it ideal for structural stylization that requires semantic awareness.
-This is fundamentally different with latent-based, solidary representation (e.g., $\mathbf{x}_{\tau}$) and pixel-level representation (e.g., ControlNet, IP-Adapter), which are uni-resolution, at single SNR level, and semantically-poor.
-Consequently, as evidenced in our experiment, these methods fail to produce meaningful structure styles and are prone to deterioration of original semantics, both in terms of structure and appearance.
-We implement the content control operator $\circ$ as injection (i.e., $a\circ b=a$).
-Semantic-aware Appearance Representation.
-Akin to structure, the appearance-related style also must also be harmoniously aligned with the image semantics.
-However, until now, the appearance delegation denoise from a random noise and hence has no interaction with the source image.
-To solve this, we propose Structure-to-Appearance injection (S2A) to propagate the high-level semantics into the generation of appearance to avoid spatial misalignment.
-Specifically, we inject self-attention value $V^{s}$ from early layers of $\epsilon^{s}_{\theta}$ to $\epsilon^{a}_{\theta}$.
-We represent the appearance of the target style as self-attention maps from all layers in the $\epsilon^{s}_{\theta}$.
+Formally, we denote the hidden feature of a ResBlock as $h_{i}(\mathbf{x}_{t})$, where $i\in\{1,2,\ldots,N_{res}\}$ indexes the ResBlocks up to $N_{res}$, with increasing spatial resolution.
+Stacking such feature from all layers forms a pyramidal structure representation of $\mathbf{x}_{0}$ at $t$:
+(4) $\kappa^{s}_{t}$ $=\{h_{i}(\mathbf{x_{t}})\}_{i\in S_{res}},$
+$\text{where }h_{i}$ $\text{ extracted from }\epsilon^{s}_{\theta}(\mathbf{x}_{t},t;\varnothing),$
+$\text{and }S_{res}$ $\subseteq\{1,2,\ldots,N_{res}\}.$
+Our representation is distinct as it captures multi-scale semantics and provides continuous guidance across the full denoising trajectory ($t\in[0,T]$).
+This fundamentally differs from methods relying on solitary control points (e.g., $\mathbf{x}_{\tau}$) or single-scale conditions (e.g., ControlNet, IP-Adapter).
+Such approaches are constrained to a fixed resolution and/or SNR, which architecturally limits their ability to generate complex structural styles.
+As evidenced in Sec.
+5, this constraint often leads to undesirable semantic trade-offs.
+With this pyramidal representation, we implement the structure control operator $\circ$ as injection (i.e., $a\circ b=a$).
+Semantic-aware Appearance Representation $\psi$.
+We represent the appearance of the target style as self-attention maps extracted from all layers of $\epsilon^{s}_{\theta}$.
+For the style to be applied harmoniously, its generation must be guided by the image’s semantics.
+However, until now, we denoise appearance delegation from Gaussian noise and hence has no information-sharing with the source image.
+To compensate for this, we propose Structure-to-Appearance injection (S2A) that propagates the high-level semantics into appearance generation.
+Specifically, we inject the self-attention value $V$ from early layers of $\epsilon^{s}_{\theta}$ to $\epsilon^{a}_{\theta}$.
 Let $N_{attn}$ denote the total number of attention blocks within the U-Net decoder, $S_{s2a}\subseteq\{1,2,\ldots,N_{attn}\}$ be the selected blocks for S2A injection.
-The style representation is:
-(5) $\sigma^{a}_{t}:=\{A^{s}_{i}\}_{i\in[1,2,\cdots,N_{attn}]}$
-[…]
-[…]
+The appearance representation at $t$ is:
+(5) $\psi^{a}_{t}=\{A^{a}_{i}\}_{i=1}^{N_{attn}},\quad\text{where }$ $A^{a}_{i}\text{ is extracted from }\epsilon^{a}_{\theta}\left(\mathbf{x}_{t},t;y,\{V^{s}_{j}\}_{j\in S_{s2a}}\right),$
+with $\{V^{s}_{j}\}_{j\in S_{s2a}}\text{ extracted from }\epsilon^{s}_{\theta}(\mathbf{x}_{t},t;\varnothing).$
 Figure 3.
 ResBlock feature map visualization.
-We apply t-sne to visualize the feature map of different feature maps in U-Net.
+We apply t-sne to visualize the feature map of different feature maps in U-Net decoder.
 The hidden features $h(\mathbf{x})$ better preserves the semantics than the ResNet feature $f(\mathbf{x})$ throughout all $T$.
@@
 Stylized images generated by DiffArtist produce high-fidelity structural and appearance-level style with semantic integrity.
-We suggest readers for more visualizations in Appendix.
-A.
+We suggest readers for more visualizations in Appendix A.
 Inspired by StyleAligned (Hertz et al., 2024) we design the appearance-style control operator $\diamond$ as the AdaIN (Huang and Belongie, 2017),
-(6) $a\diamond b=\operatorname{Var}(b)(\frac{a-\operatorname{Mean}(a)}{\operatorname{Var}(a)})+\operatorname{Var}(b).$
-The entire control process is visualized in Fig.
+(6) $a\diamond b=\sigma(b)\left(\frac{a-\mu(a)}{\sigma(a)}\right)+\mu(b).$
+We visualize the feature interactions in the right part of Fig.
 2.
-Adjusting the control layer $S_{res}$ and style strength in $\epsilon_{t}^{s}$ enables disentangled control for structure and appearance, respectively.
-3.6. MLLMs as Stylization Evaluator
-Recent research demonstrates a strong semantic-level multimodal understanding of MLLMs (Zhang et al., 2024; Kojima et al., 2022; Zhang et al., 2023c).
-Inspired by these findings, we leverage state-of-the-art MLLMs as strong zero-shot artistic evaluators.
-Due to the inherent subjectivity in artistic evaluation, we argue that obtaining an absolute score for structure or appearance fidelity can be challenging.
-Consequently, we evaluate them in a relative way, where the results from all compared methods are visible to MLLM for ranking.
-Besides, we aim to measure the fidelity, instead of the more subjective visual-appeal [^1] Note that, visual-appeal is also not equivalent to the concept of aesthetics.
-(Gombrich and Gombrich, 1995).
+Adjusting the control layer in $S_{res}$ and style strength of $\epsilon_{t}^{a}$ enables disentangled control for structure and appearance, respectively.
+3.6. Structure and Appearance Evaluation via MLLMs
+Recent research demonstrates the powerful semantic-level multimodal understanding of MLLMs (Zhang et al., 2024; Kojima et al., 2022; Zhang et al., 2023c; Li et al., 2025a, b).
+We leverage state-of-the-art MLLMs as a zero-shot evaluator to assess two key axes of our method: structure preservation and appearance fidelity.
+Crucially, our goal is to measure fidelity of stylized images, not subjective qualities like visual appeal.
+[^1] Note that visual-appeal is also not equivalent to aesthetic quality (Gombrich and Gombrich, 1995; Jiang and Chen, 2025).
 […]
@@
 Implementation Details.
-The pretrained Stable Diffusion 2.1  [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1 is used in our main experiment.
+Our experiments are built upon the publicly available Stable Diffusion 2.1 model [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1.
 We perform DDIM sampling with $T=50$ steps.
-During the inversion, the intermediate noise predictions are recorded, which are used to overwrite the input of denoiser of the content delegation at every steps.
-Further implementation details can be found in Appendix B.
-We default $S_{\text{res}}$ to be the first four ResNet layers ([1, 2, 3, 4]), and $S_{\text{s2a}}$ as the first two attention layer features ([1, 2]), unless otherwise specified.
+During the inversion, we record the intermediate noise predictions to overwrite the input of $\epsilon^{s}_{\theta}$ during denoising.
+Further implementation details are available in Appendix B.
+Experiments were conducted with a single RTX 4090-D GPU, with an approximate runtime of 2 seconds for inversion and 8 seconds for the final stylization.
+Default Parameters.
+In main experiment, we default $S_{\text{res}}$ to be the first four ResNet layers ([1, 2, 3, 4]), and $S_{\text{s2a}}$ as the first two attention layer features ([1, 2]).
 The classifier-free guidance (CFG) scale is set as $7.5$.
-Note that the default parameter corresponds to only moderate structural variations, used to set a fair comparison with other works to avoid per-image tuning.
-However, it should be noted that this paper focuses on improved controllability to allow for customization, allowing user to personalize according to their taste.
-Experiments are conducted with a single RTX 4090-D GPU.
-The inversion process takes approximately 2 seconds, and the subsequent stylization process takes around 8 seconds.
+These default control parameters correspond to moderate structural and appearance variations, used to set a fair comparison with existing works to avoid per-image parameter tuning.
+However, it should be noted that users can adjust these parameters for customization.
 Compared Methods.
 We compare our method against existing text-driven stylization and manipulation methods: DDIM Inversion (Song et al., 2020), CLIPStyler (optimization-based) (Kwon and Ye, 2022), DiffStyler (Huang et al., 2024), Plug-and-Play (PnP), Prompt2Prompt (with null text inversion) (P2P)  (Mokady et al., 2023; Hertz et al., 2022), ControlNet (Zhang et al., 2023a), and InstructPix2Pix (Brooks et al., 2023).
-Additionally we consider a baseline named SD, which generate image with the Stable Diffusion according to $y$.
-We also indirectly compare our method with reference-based stylization methods, including AdaIN (Huang and Belongie, 2017), ArtFlow (An et al., 2021), StyTr2 (Deng et al., 2022), StyleID (Chung et al., 2024), StyleAligned (Hertz et al., 2024) (with ControlNet) and InstantStyle (Wang et al., 2024a) (with ControlNet).
+Additionally, we consider a baseline named SD, which generates images with Stable Diffusion according to $y$.
+We also indirectly compare our method with reference-based stylization methods, including AdaIN (Huang and Belongie, 2017), ArtFlow (An et al., 2021), StyTr2 (Deng et al., 2022), StyleID (Chung et al., 2024), StyleAligned (Hertz et al., 2024) (with ControlNet), and InstantStyle (Wang et al., 2024a) (with ControlNet).
 Images generated by SD are used as reference.
-Conventional Metrics: LPIPS (Zhang et al., 2018) measures the content preservation by calculating the perceptual distance from source and stylized image.
-For style fidelity, we leverage CLIP Score (Radford et al., 2021) and Pick Score (Kirstain et al., 2023), both quantify the alignment between the stylized image $\hat{I}$ and prompt $y$.
-We also include a human study crowd-sourced from $n_{1}=200$ users and report the average preference rate.
-MLLM-based Metrics: We rank the stylization fidelity in ascending order, and normalize the rankings across all test samples.
-Therefore, a score closer to 1 indicate stronger fidelity.
-We use Gemini-v2.0-flash is used due to its strong multimodal task-solving capability and empirically superior alignment with human.
+Conventional Metrics: LPIPS (Zhang et al., 2018) measures the content preservation by calculating the feature distance between the source and stylized image.
+For style fidelity, we leverage CLIP Score (Radford et al., 2021) and Pick Score (Kirstain et al., 2023), both of which quantify the alignment between the stylized image $\hat{I}$ and prompt $y$.
+We also include a human study crowd-sourced from $n_{1}=200$ users and report the average preference rate for our method.
+MLLM-based Metrics: We prompt the MLLM to rank the fidelity of $k$ stylized images from best (rank 1) to worst (rank $k$).
+We normalize the integer ranking and average it over the whole evaluation set.
+Therefore, a score closer to $1$ indicates a stronger fidelity.
+We use Gemini-v2.0-flash for its strong multimodal capability.
 The full prompt templates can be found in Appendix C.
@@
 Qualitative Comparisons.
-We first provide a comprehensive comparison against previous text and reference-based stylization methods, visualized in Fig.
+We first provide a comprehensive comparison against previous methods, visualized in Fig.
 4-(a, b).
-(a): When compared with text-driven methods, DiffArtist is the best at following the style prompt while maintaining content integrity.
+(a): Compared with text-driven methods, DiffArtist is the best at following the style prompt while maintaining semantic integrity.
 Our method enables harmonious structural variations, such as pixelation, without compromising intricate details like facial identity and hair.
-In contrast, the compared methods (e.g., DiffStyler, Plug-and-Play) may introduce undesired modifications that violates semantics or produce misaligned styles (e.g., InstructP2P, ControlNet).
-(b): When broadly compared with reference-based methods, our method still stands out for its capability to curate both appearance and structural variations.
-To fully demonstrate the results, we suggest readers for additional visualizations in Appendix A.
+By contrast, the compared methods may produce misaligned styles (e.g., CLIPStyler, Plug-and-Play) or introduce undesired modifications that violate semantics (e.g., DiffStyler, ControlNet).
+(b): When broadly compared with reference-based methods, DiffArtist still stands out for its high stylization fidelity from two perspectives.
+To fully demonstrate the superiority, we highly suggest readers for additional visualizations in Appendix A.
 Quantitative Comparison.
-For our quantitative evaluation, we first sample 50 art styles from WikiART, with both abstract (e.g., “Cubism”) and realistic styles (e.g, “High-renaissance”), which are diversified by GPT-4o in terms of description.
-This sets a wide spectrum of styles to align with real-world applications.
+For our quantitative evaluation, we first sample 50 art styles from WikiART, with both abstract (e.g., “Cubism”) and realistic styles (e.g., “High Renaissance”), which are further diversified by GPT-4o in terms of description.
+This diversification sets a broad spectrum of styles to align with real-world user inputs.
 The content comprises 50 images from MSCOCO (Lin et al., 2014) and 50 photorealistic images generated by another model (Ideogram, 2024).
-For each content image we randomly draw $10$ style prompts from all possible styles, resulting in a total of $1,000$ combinations for comparison.
-The comparative results are presented in Tab.
-1.
+For each of the $100$ content images, we randomly draw $10$ style prompts from all possible styles, resulting in a total of $1,000$ unique combinations for comparison.
+Tab.
+1 presents the results.
 For conventional metrics, DiffArtist achieves an LPIPS of 0.52, a CLIP Score of 25.91, and a PickScore of 20.51, outperforming most of the compared methods.
-However, these metrics discard either the content or style reference, and they do not explain stylization quality in terms of structure and appearance.
-For example, the baseline SD has be best CLIP Score and PickScore but it completely discard the content information in the original image.
+However, these metrics do not measure stylization quality in structure and appearance.
+As a simple counter-example, the baseline SD has the highest CLIP Score and PickScore, whereas it is not even performing stylization.
 We include these scores solely for reference.
-When evaluated with MLLM-based metrics, DiffArtist attains the highest average score.
+When evaluated with MLLM-based metrics, DiffArtist attains the highest average score of 0.64.
 Specifically, our method achieves the second-highest structure score of 0.61, demonstrating DiffArtist’s effectiveness in generating structural styles.
-The only stronger baseline is P2P, which is designed for image editing, with relatively weak stylization capability.
-Besides, our method achieves the best appearance score, showcasing the advanced capability of DiffArtist in capturing high-fidelity appearance details embedded in style prompt.
-In human evaluations, our method is preferred 67.8% of users, further validating its superiority.
+While the editing-focused P2P method scores higher, it does so by sacrificing stylistic strength, evidenced in qualitative comparisons.
+Besides, our method achieves the best appearance fidelity score, confirming its superior ability to render the appearance details from the text prompt.
+In human evaluations, our method is preferred by an average of 67.8% of users in pairwise comparison, further validating its superiority.
 5. Analysis and Discussion
-This section provided extended comparisons and discussions.
+This section provide in-depth analysis on the proposed system.
 In Sec.
-5.1 we provide an in-depth comparison of different style control methods, both qualitative and quantitatively.
+5.1, we analyze the controllability of DiffArtist in detail.
 Sec.
-5.2 validate the effectiveness of MLLMs as style evaluator.
+5.2 validates the effectiveness of MLLMs as style evaluators.
 Sec.
-5.3 we provide ablations on delegation and S2A injection.
-We conclude this section with a discussion in Sec.
+5.3 provides ablations on delegations and the S2A injection.
+We conclude with a discussion in Sec.
 5.4.
 Figure 5.
-Comparisons between representative style control methods.
-DiffArtist offers significantly stronger controllability than other stylization methods.
-(Left): DiffArtist allows smooth and artistically meaningful structural stylization strength control, while maintaining appearance style strength.
+DiffArtist offers disentangled and stronger controllability.
+(Left): DiffArtist enables smooth and artistically meaningful structural stylization at varying degree, without violating appearance style strength.
 (Right): DiffArtist allows fine-grained control of appearance-related style strength while preserving structural and semantic integrity.
-5.1. DiffArtist has Fine-grained Controllability
-This section aims to validate the unprecedented dual controllability of DiffArtist.
-Specifically, we provide fine-grained comparisons of the following three groups of representative methods, configured with different control levels: (a) Semantic pyramid: include DiffArtist and DiffArtist implemented with Plug-and-Play structure representation, $f(\mathbf{x})$ (Ours + PnP), and Ctrl-X; (b) Pixel-level map: include InstantStyle (Cui et al., 2023), which is based on ControlNet (Zhang et al., 2023a) to use pixel-level maps as additional conditioning; (c) Noise inversion, which corresponds to the DDIM baseline.
-For content control in group (a), we set 5 levels, using the following control layers: ($\varnothing$, [1],[1-4],[1-6],[1-8]); for group (b), we evenly sample their respective control strength hyperparameters; and we sample $\tau=[0,5,10,15,20]$ for group (c).
-Style strength is controlled by sampling CFG weights in [2.5,5,7.5,10] for all groups except for Ctrl-X, which is achieved by adjusting its appearance schedule parameter.
-Qualitative Comparison.
+5.1. DiffArtist has Strong Controllability
+This subsection validates the unprecedented dual controllability of DiffArtist.
+To achieve this, we conduct a fine-grained comparative analysis against representative methods, categorizing them by their core control mechanism: (a) Semantic pyramid: include DiffArtist, DiffArtist implemented with Plug-and-Play structure representation, $f(\mathbf{x})$ (Ours + PnP), and Ctrl-X; (b) Pixel-level map: include InstantStyle (Cui et al., 2023), which is based on ControlNet (Zhang et al., 2023a); (c) Noise inversion, which corresponds to the DDIM baseline.
+For structure control, we define five levels from weakest to strongest.
+For group (a) we use the following control layers: ($\varnothing$, [1],[1-4],[1-6],[1-8]); for group (b), we evenly sample their respective control strength parameters; and we use $\tau=[0,5,10,15,20]$ for group (c).
+Appearance strength is controlled by sampling CFG weights in [2.5,5,7.5,10] for all groups except for Ctrl-X, which is achieved by adjusting its appearance schedule parameter.
+Qualitative Comparison on Control.
 As visualized in Fig.
-5, we observe the strength of structure and appearance style are the most adjustable in DiffArtist, with harmonious, consistent, and disentangled interpolations across a sequence of control levels.
-Specifically, the compared methods (Ours+PnP) and group (b) primarily produce style at the appearance level, failing to capture the essential fragmented shape of Cubism at varying levels of abstraction.
-When tuned with different appearance strength, these methods are also incapable of producing artistically meaningful interpolations.
-While DiffArtist produce gradually bolder strokes, the compared methods only gives more saturated results.
-This experiment also demonstrate the effectiveness of content representation: the pixel-map based representation can severely violate the semantic integrity, like the example of InstantStyle where the face is barely recognizable.
-The inversion-based representation, on the other hand, entangles structure and appearance, making it difficult to control one without affecting the other.
-Further visualization in Appendix.
-A more clearly demonstrate the strength of DiffArtist.
-Quantitative Comparisons.
-We provide quantitative experiments to further validate the observations from the visualization above.
-We consider two properties of control: quality and editability.
-(1) We compare the quality, measured as structure and appearance fidelity scores at different control levels, across different methods.
+5, DiffArtist demonstrates superior controllability, with harmonious, consistent, and disentangled interpolations across a sequence of control levels.
+It correctly captures the essential geometric principles of Cubism, whereas competing methods merely apply a superficial texture (e.g., Ours+PnP) or fail to produce meaningful structural variations (e.g., InstantStyle).
+This nuanced control is further evident when adjusting style strength; DiffArtist provides an artistically meaningful interpretation by producing bolder strokes according to style prompt, while other methods resort to simplistic and often undesirable increases in color saturation.
+Most critically, our approach preserves semantic integrity.
+It avoids the catastrophic failures of pixel-map methods like InstantStyle, which can render the face unrecognizable, and also prevents the facial structure corruption seen in inversion-based methods that inherently entangle form with appearance.
+The extended visualizations in Appendix A with diverse style and content further confirm the superiority of DiffArtist.
+Quantitative Comparisons on Control.
+We provide quantitative experiments to substantiate our visual observations, evaluating control based on two properties: fidelity and editability.
+We measure the fidelity as structure and appearance MLLM scores at different control levels.
 The results are reported in Tab.
 2.
-DiffArtist outperforms others significantly and consistently, demonstrating superior control quality.
-We further consider (2) the editability and disentanglement properties of the control within each method.
-For instance, when increasing structure control, one would expect a gradually stronger structure style, covering a wide range (editability) with minimal influence on the appearance-related style (disentanglement), and vice versa.
-In other words, during a sequence of control, the structure fidelity should exhibit monotonic growth while the appearance score should resemble a random distribution.
-To quantify this intuition, we use Spearman’s coefficient of correlation $\rho$ to measure the monotonicity of control, and Kendall’s coefficient of concordance $W$ to quantify the randomness of sequential manipulation.
-A $\rho$ close to 1 indicates stronger monotonic correlation, while a $W$ close to 0 indicates random ranks.
+DiffArtist outperforms others significantly and consistently, demonstrating superior control fidelity.
+The editability defines the quality of the manipulation itself, which we assess via three criteria: range, monotonicity, and disentanglement.
+An ideal control of stylization should cover a wide range of stylistic effects (a large spread in MLLM scores), exhibit predictable monotonicity (Spearman’s $\rho\approx 1$ for the target attribute), and maintain disentanglement from other attributes.
+We measure the disentanglement using Kendall’s $W$ on the off-target scores, where a stable, unaffected score sequence yields $W\approx 0$.
+For instance, when controlling structural strength, $W$ for the appearance score is desired to be near 0.
 As shown in Tab.
-2, the control of DiffArtist is the most editable and disentangled, reaffirming the observation of the visualizations in Fig.
+2, the control of DiffArtist is the most editable: it covers the broadest range of effects, demonstrates the strongest monotonicity, and achieves the best disentanglement, reaffirming the superior control visualized in Fig.
 5.
@@
 Fidelity of structure and appearance control via cross-method comparison.
-The values corresponds to structure and appearance score (MLLM, $\uparrow$), respectively.
-Note that the magnitude of scores are only comparable within each column.
+The values correspond to the structure or appearance score (MLLM, $\uparrow$).
+Note that the magnitudes of scores are only comparable within each column.
 The best result for each column is in bold.
@@
 Control editability and disentanglement via inter-method comparison.
-Higher $\rho$ indicates stronger correlation while lower $W$ means ranks are random.
+Higher $\rho$ indicates stronger monotonicity while lower $W$ means ranks are indistinguishable.
 When controlling from one perspective, a high $\rho$ is desired for editability, and a low $W$ for the other aspect is expected for disentangled control.
@@
 $W(A)\downarrow$ 0.32 0.44 0.45 0.34 0.72 $\rho(A)\uparrow$ 0.71 0.42 0.35 0.26 0.68
-5.2. MLLMs are Human-Aligned Style Transfer Metrics
-We evaluate how each stylization metric is aligned with human perception by calculating the statistical correlation with rankings from human feedback.
-To achieve this, we first construct a comparison set of 800 stylized images [^3] The images tested here do not overlap with the experiment in Tab.
+5.2. MLLMs are Human-Aligned Stylization Evaluators
+We evaluate how each stylization metric aligns with human feedback by calculating the statistical correlation with the rankings from human feedback.
+To achieve this, we first construct a comparison set of 800 stylized images [^3] The images evaluated here do not overlap with the main experiment in Tab.
 1, and compare how human and MLLM preferences correlate.
-For human feedback, we consider two groups of users.
-For the non-expert group, we distributed questionnaires on a crowd-sourced platform, where each participant was required to complete randomly sampled ranking tasks.
-A total of $n_{1}=200$ participants took part in this study.
-Filters were applied to validate and clean the raw feedback.
-We also recruited an expert group of $n_{2}=12$ participants (including the author(s)), who have experience in painting or creative design.
-We utilize spearman’s $\rho$ to measure the alignment between metrics and human feedback.
-The averaged (of all content-style pairs) coefficient $\rho$ for both groups is reported in Tab.
+To gather human feedback, we consider two groups of users.
+For the non-expert group, we recruited a large-scale group of $n_{1}=200$ participants through a crowdsourcing platform.
+Each participant performed a series of randomly sampled ranking tasks.
+To ensure the integrity of the collected data, we implemented attention checks and consistency filters to remove unreliable responses.
+We also recruited an expert group of $n_{2}=12$ participants with knowledge of fine art.
+We measured the alignment between each metric’s rankings and the human-derived preferences using Spearman’s rank correlation ($\rho$).
+The averaged (of all content-style pairs) $\rho$ for both groups is reported in Tab.
 4.
-As the table shows, the MLLM-based metrics are better aligned with human perception, suggesting its effectiveness as an evaluation metric for style transfer.
-Table 4. Metrics correlation with human feedback. We report correlation $\rho$ and combined significance $p$. The VLM-metrics align better with human evaluation for both expert and non-expert users.
+As the table shows, the MLLM-based metrics are better aligned with human perception, validating its effectiveness as an evaluation metric for style transfer.
+Table 4. Metrics correlation with human feedback. We report correlation $\rho$ and combined significance pp. The MLLM scores show stronger alignment with both expert and non-expert users.
 Corr.
 (Non-expert) Corr.
-(Expert) Metrics $\rho$uparrow p-value $\downarrow$ $\rho$uparrow p-value $\downarrow$ S SSIM (Wang et al., 2004) 0.29 0.12 0.25 0.14 MLLM (GPT-4o) 0.44 0.004 0.34 0.20 MLLM (Gemini v2) 0.42 0.02 0.45 0.03 A CLIP Score 0.05 0.73 0.01 0.75 Pick Score (Kirstain et al., 2023) 0.27 0.11 0.25 0.13 MLLM (GPT-4o) 0.25 0.05 0.22 0.06 MLLM (Gemini v2) 0.48 0.04 0.41 0.02
+(Expert) Metrics $\rho$ $\uparrow$ p-value $\downarrow$ $\rho$ $\uparrow$ p-value $\downarrow$ S SSIM (Wang et al., 2004) 0.29 0.12 0.25 0.14 MLLM (GPT-4o) 0.44 0.004 0.34 0.20 MLLM (Gemini 2.0) 0.42 0.02 0.45 0.03 A CLIP Score 0.05 0.73 0.01 0.75 Pick Score (Kirstain et al., 2023) 0.27 0.11 0.25 0.13 MLLM (GPT-4o) 0.25 0.05 0.22 0.06 MLLM (Gemini 2.0) 0.48 0.04 0.41 0.02
 5.3. Ablations
 The delegations enable dual controllability.
-The control mechanism of DiffArtist is built upon the two delegations.
-Removing either one will damage the dual controllability.
-Therefore we compare removing the delegations from full method with the default parameters and summarize the result in Tab.
-5.
-The full methods achieves the best results, demonstrating the synergistic effect of delegations.
-Table 5.
-Ablation on delegation branches.
-Content delegation reduces unrelated content modification, while the style delegation boosts stylization strength.
-Controls from both delegations are disentangled from each other.
+DiffArtist’s controllability stems from delegating structure and appearance generation to separate processes.
+To test the necessity of each, we created two ablated variants for comparison where the structure or appearance delegation is removed.
+Tab.
+5 presents the result of this ablation.
+The full methods achieves the best results, demonstrating the synergistic effect of delegations for dual controllability.
+Table 5. Ablation on delegation branches. The proposed two delegations are complementary to each other, and the full methods achieves the highest fidelity.
 Metric Method full w/o structure w/o appearance LPIPS $\downarrow$ 0.51 0.76 0.42 CLIP Score $\uparrow$ 25.91 27.69 21.75 Pick Score (Kirstain et al., 2023) $\uparrow$ 20.55 20.57 20.41 Structure (MLLM) $\uparrow$ 0.72 0.37 0.33 Appearance (MLLM) $\uparrow$ 0.62 0.59 0.22
-S2A injection promotes uniform spatial distribution of style strength in the style delegation process, thereby avoiding artifacts in the final stylization result.
+S2A injection promotes semantic-aligned spatial distribution of style strength in the style delegation process, thereby avoiding artifacts in the final stylization result.
 We visualize the denoised style image (from appearance delegation) and the final stylization result in Fig.
 6.
-As observed in the examples, without the S2A injection, the appearance delegation may generate irrelevant structures with uneven appearance distribution, adversely affecting the corresponding regions of the stylization results.
-Figure 6. Ablation on S2A injection. S2A injection propagate the high-level semantic to the appearance generation. It avoids spatial-misalignment of appearance-style strength.
+Without S2A injection, the appearance delegation fails to align with content semantics, generating an appearance reference image with undesired patterns and an uneven texture distribution.
+These flaws manifest directly in the final output as distracting visual artifacts.
+In contrast, the full model leverages S2A to produce a coherent style representation, resulting in a clean and high-quality final image.
+Figure 6. Ablation on S2A injection. S2A injection propagates the high-level semantic to the appearance generation. It avoids spatial misalignment of appearance-style strength.
 5.4. Limitations & Future work
-Disentangling the the interplay of structure and appearance in artistic styles is a long-standing challenge.
-Despite the progress we made in this paper, significant work remains be done.
+While DiffArtist marks a significant step towards disentangled structure and appearance control, we identify several limitations that open exciting avenues for future research.
 For instance, the structure control in DiffArtist is at a global level, and it cannot control the structure for each object separately.
 Many art styles exhibit mixed structure variation, like Surrealism and Collage art.
-We believe that the development of dense structure evaluators with a 2D reward signal is a promising direction (Liang et al., 2024), which would require a deeper understanding of both art and computation.
+Developing dense structure evaluators with 2D feedback signals is a promising direction (Liang et al., 2024), which may be further utilized as a reward model for reinforcement learning (Black et al., 2023; Clark et al., 2023).
 6. Conclusion
 We present the first exploration of structure- and appearance-controllable image stylization.
-Our contributions include DiffArtist, a generator that fully disentangles structure and appearance during the diffusion process, and a human-aligned evaluator to assess structural and appearance fidelity at the semantic level.
-Our extensive analysis prove that semantically rich representations are essential for both structure and appearance, allowing for high-fidelity and editability during control, similar to that of a human artist.
-We hope these design principles broadly inspire future art generation methods to produce truly artistically meaningful paintings.
+Our contributions include DiffArtist, a styler that fully disentangles structure and appearance during the diffusion process, and a human-aligned evaluator to assess structural and appearance fidelity at the semantic level.
+Our extensive analysis proves that semantically-rich representations are essential for both structure and appearance style.
+We demonstrated that our design allows for high style fidelity and controllability, similar to that of a human artist.
+We believe the objective established in this paper—to stylize in both structure and appearance—offer a roadmap for the next generation of generative art tools to produce artistically meaningful paintings.
+7. Acknowledgment
+This research was supported by the Hong Kong Research Grants Council (GRF-15229423).
 Appendix A Additional Qualitative Results
@@
 Tradeoff between Structure and Appearance Style Control.
-We present the experimental (solid lines) and simulation (dotted lines) trends of structural and appearance variation in diffusion process.
+We present the experimental (solid lines) and simulation (dotted lines) trends of structural and appearance variation in the diffusion process.
 Due to quadratic growth, highly noisy steps are required to achieve strong appearance styles, which are associated with significant structural variation, which can violate semantics.
 Top: Example stylization results starting from different steps, using the prompt: “Fauvism painting”.
-When the appearance strength is high ($t=0.8T$), the structure (legs of the horse) are incorrectly modified.
+When the appearance strength is high ($t=0.8T$), the structure (legs of the horse) is incorrectly modified.
 Figure 8.
@@
 Visualization Additional appearance stylization is in Fig.
-10, their source image is in Fig.
+10, the source image is in Fig.
 11.
-A grid of different image with different style is in Fig.
+A grid of different images with different styles is in Fig.
 12.
@@
 Additional qualitative comparisons on stylization.
-We show extended comparison with previous reference-based method in Fig.
+We show an extended comparison with the previous reference-based method in Fig.
 15.
-More qualitative comparisons with existing text-driven image stylization and manipulation methods could be found in Fig.
+More qualitative comparisons with existing text-driven image stylization and manipulation methods can be found in Fig.
 16 and Fig.
@@
 19.
-These results clearly demonstrate the advantage of DiffArtist in providing disentangled structural and appearance-level style control.
-In particular, the Ctrl-X, as an image editing method, produce less visually pleasing result when applied to image stylization.
-This is because they have a different definition of appearance and structure for the editing of real photos.
+These results demonstrate the advantage of DiffArtist in providing disentangled structural and appearance-level style control.
+In particular, the Ctrl-X, as an image editing method, produces less visually pleasing results when applied to image stylization.
+This is because they have a different definition of appearance and structure for editing real photos.
 Figure 10. Addition results for DiffArtist (with default control parameters). The image semantics are preserved with strong and high-fidelity styles harmoniously integrated.
@@
 1, we obtain:
-(7) $\mathbf{x}_{t-1}$ $=\mathcal{A}_{t}\mathbf{x}_{t}+\mathcal{B}_{t}\epsilon_{\theta}(\mathbf{x}_{t},t;y),\quad\mathcal{A}_{t}:=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}$
-$\mathcal{B}_{t}$ $:=\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t-1})}}{\sqrt{\alpha_{t}}}$
+(7) $\mathbf{x}_{t-1}$ $=\mathcal{A}_{t}\mathbf{x}_{t}+\mathcal{B}_{t}\epsilon_{\theta}(\mathbf{x}_{t},t;y),\quad\mathcal{A}_{t}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}$
+$\mathcal{B}_{t}$ $=\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t-1})}}{\sqrt{\alpha_{t}}}$
 Based on above formulation, the full stylization process could be expressed as:
@@
 8, the conceptual denoising term from $\mathbf{x}_{T}$ to $\mathbf{x}_{\tau}$ preserves the structure and appearance in $I_{c}$.
-The stylization trajectory from $\mathbf{x}_{\tau}$ to $\mathbf{x}_{t}$ is what actually computed, which introduces the desired appearance based on the prompt $y$ but may also lead to uncontrolled structure generation.
-In this paradigm, the preservation of original appearance and generation of new structure are side-effects that are undesirable.
-Intuitively, larger $\tau$ leads to a shorter trajectory of $\mathbf{x}_{T}\to\mathbf{x}_{\tau}$, resulting in a stronger appearance in $\hat{\mathbf{x}}_{0}$ with weaker structure preservation (uncontrolled structure stylization), while lower $\tau$ sacrifices stylization strength for stronger structure preservation.
+The stylization trajectory from $\mathbf{x}_{\tau}$ to $\mathbf{x}_{t}$ is what is actually computed, which introduces the desired appearance based on the prompt $y$ but may also lead to uncontrolled structure generation.
+In this paradigm, preserving original appearance and generating new structure are undesirable side effects.
+Intuitively, a larger $\tau$ leads to a shorter trajectory of $\mathbf{x}_{T}\to\mathbf{x}_{\tau}$, resulting in a stronger appearance in $\hat{\mathbf{x}}_{0}$ with weaker structure preservation (uncontrolled structure stylization), while a lower $\tau$ sacrifices stylization strength for stronger structure preservation.
 In other words, one can not arbitrarily control the strength of appearance and structure without affecting the other.
@@
 2 and Eqn.
-8, we can quantitatively assess the strength of content preservation and stylization in diffusion process under a particular noise schedule $\alpha_{1:T}$.
-Specifically, we further assume the SNR of structure and appearance has linear relationship with that of the $\mathbf{x}_{t}$ for $t$:
+8, we can quantitatively assess the strength of content preservation and stylization in the diffusion process under a particular noise schedule $\alpha_{1:T}$.
+Specifically, we further assume the SNR of structure and appearance has a linear relationship with that of the $\mathbf{x}_{t}$ for $t$:
 (9) $\operatorname{SNR}(\mathbf{z}_{t}^{c})\propto\frac{\alpha_{t}}{1-\alpha_{t}};\operatorname{SNR}(\mathbf{z}_{t}^{s})\propto\frac{\alpha_{t}}{1-\alpha_{t}}$
@@
 Specifically, we assume that the relative significance of each unweighted denoising step $\epsilon(\mathbf{x}_{t},t;y)$ on the final stylized image remains consistent across different timesteps for both structure and appearance.
-In other words, we assume the SNR of structure and appearance has linear relationship with that of the $\mathbf{x}_{t}$ at time $t$ as characterized by the noise schedule $\alpha_{1:T}$:
+In other words, we assume the SNR of structure and appearance has a linear relationship with that of the $\mathbf{x}_{t}$ at time $t$ as characterized by the noise schedule $\alpha_{1:T}$:
 (10) $\operatorname{SNR}(\mathbf{z}_{t}^{c})\propto\frac{\alpha_{t}}{1-\alpha_{t}};\operatorname{SNR}(\mathbf{z}_{t}^{s})\propto\frac{\alpha_{t}}{1-\alpha_{t}},$
 It is important to note that we do not assume that the relative proportions of structure and appearance are identical at each denoising step.
-With above assumption, the effect of varying $\tau$ on structure and appearance of final stylized image could be derived in close-form. In practice, we use the following code to calculate iteratively:
+With the above assumption, the effect of varying $\tau$ on the structure and appearance of the final stylized image could be derived in closed form.
+In practice, we use the following code to calculate iteratively:
 def cum_score(low_t, hi_t, alphas):
@@
 To address this issue, we adopt an alternative strategy by randomly sampling 500 Gaussian noise as the $\mathbf{x}_{T}$ of content, which are paired with 500 content prompts.
-We treat the images that denoised using content prompts for $\tau=T$ steps as the content image, which simulate a perfect inversion technique.
-To stylize an image, we first denoise the $\mathbf{x}_{T}$ with content prompt for $\tau$ steps, which is subsequently denoised with the style prompt for $T-\tau$ steps.
-The LPIPS between the stylized image and content image is used as the empirical structural strength, whereas CLIP Deception score (correct classification rate among a set of styles) is used as the empirical appearance strength.
-The following 10 style prompts is used:
+We treat the images denoised using content prompts for $\tau=T$ steps as the content image, which simulates a perfect inversion technique.
+To stylize an image, we first denoise the $\mathbf{x}_{T}$ with the content prompt for $\tau$ steps, which is subsequently denoised with the style prompt for $T-\tau$ steps.
+The LPIPS between the stylized image and the content image is used as the empirical structural strength.
+In contrast, the CLIP Deception score (correct classification rate among a set of styles) is used as the empirical appearance strength.
+The following 10 style prompts are used:
   • ”watercolor style”
@@
   • ”abstract expressionism style”
-The result for both simulation and empirical results are in Fig.
+The results for both simulation and empirical results are in Fig.
 7.
-The result show a good fit, and it turns out that the structure modification appears to be linear, with the stylization strength is quadratic with respect to $\tau$.
+The result shows a good fit, and it turns out that the structure modification appears to be linear, with the stylization strength being quadratic with respect to $\tau$.
 Moreover, this result further evidenced the issue of S-A entanglement in the diffusion process.
@@
 C.1. Implementation Details
-The stylized images, style prompt, and the instruction prompts are feed to MLLM for inference.
+The stylized images, style prompt, and the instruction prompts are fed to MLLM for inference.
 We compose stylized images as a grid image with numbers at the top-left corner.
@@
 If a participant answers an attention-check question incorrectly, all of their responses will be marked invalid.
-Response that are made with less than 20 seconds are also removed.
+Responses that are made with less than 20 seconds are also removed.
 Table 6. Prompt templates for MLLM-based metrics. [IMG], [STYLE] and [NUM_METHOD] is the placeholder for combined image, style and number of methods, respectively.
@@
 Empirically, our default configuration $[1,2]$ strikes an optimal balance between enhancing style detail and preserving essential content structure.
-In this section we provide additional ablations to study the effect of proposed S2A design.
+In this section, we provide additional ablations to study the effect of proposed S2A design.
 Ablation on S2A layers $S_{S2A}$.
@@
 D.3. Failure Case
-In our experiment, we identify a rare ($<$1%) and special failure case in the proposed methods.
+In our experiment, we identify a rare (<<1%) and special failure case in the proposed methods.
 Specifically, for certain content image, its stylization result will consistently contains black and white chessboard-pattern artifacts.